### v2

# Open Agent Specification (Agent Spec) Technical Report

Abstract
Open Agent Specification (Agent Spec) is a declarative language that allows AI agents and their workflows to be defined in a way that is compatible across different AI frameworks, promoting portability and interoperability within AI Agent frameworks.
Agent Spec aims to resolve the challenges of fragmented agent development by providing a common unified specification that allows AI agents to be designed once and deployed across various frameworks, improving interoperability and reusability, and reducing redundant development efforts. Additionally, Agent Spec facilitates development tools and portability, allowing AI agents to be defined independently of their execution environment and enabling teams to exchange solutions without implementation-specific limitations.
Agent Spec benefits four key groups: (i) Agent developers, who gain access to a superset of reusable components and design patterns, enabling them to leverage a broader range of functionalities; (ii) Agent framework and tool developers, who can use Agent Spec as an interchange format and therefore benefit from the support of other frameworks as well as other tools; (iii) Researchers, who can achieve reproducible results and comparability, facilitating more reliable and consistent outcomes; (iv) Enterprises, which benefit from faster prototype-to-deployment, increased productivity, as well as greater scalability and maintainability for their AI agent solutions. This technical report provides an overview of the technical foundations of Agent Spec, including motivation, benefits, and future developments.
1 Introduction
This technical report introduces Agent Spec, a declarative, framework-agnostic configuration language designed to define AI agents and their workflows with high fidelity. By providing a unified representation, Agent Spec enables seamless portability and interoperability of agents across diverse frameworks, ensuring consistent behavior and execution across various AI Agent frameworks. Agent Spec serves as an abstraction layer above framework-specific specifications, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints.
1.1 Background & Motivation
Existing agentic frameworks have their own strengths and limitations. While flexibility is crucial to accommodating diverse agentic design patterns, currently available frameworks do not fully support all of them. Table 1 outlines key objectives a comprehensive framework should cover and describes how Agent Spec addresses each of these requirements.
Table 1: Challenges and objectives in agentic frameworks and solutions that we define with Agent Spec.
Topic Objective Open Agent Specification
Agents
Author and edit conversational solutions
such as ReAct [12] style agents.
Ability to define components that represent "resources"
(with no impact on conversation flow routing) such as
LLMs, memory, prompt templates, tools, other Agents
which can be attached to other components’ properties.
Flows
Author and edit graph-style workflows
such as multi-step business processes.
Ability to define components that have a well-defined
execution path such as a directed graph which may
contain conditional routes or loops.
Control flow routing
Ability to define the flow of execution
between components in the system,
including conditional or looping flows.
Explicit "relationships" (edges) for control flow
supporting one-to-many and many-to-one transitions. The
specific transition to be taken is determined at
runtime for a given execution.
I/O routing
Ability to define which outputs are
consumed by which inputs for fixed
flow components. This includes
conditional or looping flows, and outputs
that are consumed in multiple places.
Explicit "relationships" (edges) for data flow. Support
for one-to-many and many-to-one transitions where the specific
transition to be taken is determined at runtime for a
given execution.
Nesting/Composition
Reuse components or solutions at any
level of granularity in other solutions.
Includes multi-agent composition.
Definition of components that can be encapsulated and
reused elsewhere. Established required properties for
such components. Varying levels of visibility of
exchanged messages between components.
Orchestration of execution order and dependency either
statically or dynamically.
Referencing
Ability to define a component instance
once and refer to it in multiple places.
"Reference" syntax to specify property values by
referencing and reuse. Provision of components for
defining an element and for using them.
For example, an Agent component to define agent and
an AgentNode in a Flow to use it in a certain place,
or a Tool component to define a tool and a ToolNode
in a Flow to use the tool in a specific place.
Moreover, different frameworks are usually parameterized and configured by different means. This makes porting solutions between frameworks a tedious and error-prone process, hindering collaboration and knowledge sharing among teams developing agentic applications. Consequently, a declarative, portable, and framework-agnostic representation of agentic designs is needed.
For instance, ONNX [8] has revolutionized deep learning by providing a consistent way to port ML models across different frameworks (e.g., PyTorch [9], TensorFlow [11]). Similarly, Agent Spec aims to establish a unified representation for AI agents, enabling seamless interoperability and execution across diverse agentic frameworks. Just as ONNX allows models to be trained in one framework and executed in another, Agent Spec allows AI agents to be designed once and deployed across multiple platforms without reimplementation.
1.2 Objectives of Open Agent Specification
Agent Spec aims to define a common representation and act as a superset of capabilities for AI agents and workflows. Agent Spec graphs can be executed in different frameworks by means of runtimes that map Agent Spec components to the underlying framework’s primitives. Agent Spec, along with importers and exporters for various underlying frameworks, enables transparent portability of agent workflows. Agent Spec also provides methods to validate agent workflows before execution, ensuring they are syntactically and logically correct and executable.
1.3 Key Benefits of Agent Spec
1.3.1 Portability & Reusability
  • Framework-Agnostic: Agent Spec abstracts agent definitions from specific implementations by providing a unified declarative representation, allowing agents to operate uniformly across multiple platforms (e.g., AutoGen, LangGraph, OCI Agents) without any need for reimplementation.
  • Modular Design: Agent Spec’s component-based structure fosters reusability and extensibility, enabling enterprises to build complex agentic systems by assembling and configuring standardized components.
  • Support for Complex Agentic Flows: Agent Spec facilitates the creation of intricate agentic workflows by supporting various modular components and reusable Agent Spec-defined patterns, enabling streamlined development of advanced AI agent systems.
  • Portable Execution Graphs: Agent Spec enables execution graphs to be described in a way that is independent of the underlying runtime implementation, ensuring seamless portability across frameworks. Converters between the representations of other frameworks to Agent Spec would simplify the porting of previously developed agents to Agent Spec-compatible frameworks.
1.3.2 Interoperability & Compatibility
  • SDK Support: Agent Spec provides SDKs in various programming languages (starting with Python), supporting serialization and deserialization of agents into Agent Spec configurations. This simplifies development, deployment, and debugging across different environments.
  • Robustness & Consistency: Agents defined in Agent Spec can be built on top of components that have been defined and validated within the Agent Spec specification, leading to higher reliability and predictability.
  • Define Once, Run on Multiple Frameworks: Agentic systems leveraging Agent Spec can run on various frameworks through runtime adapters. Agent Spec’s open specification allows developers to implement these adapters to support any framework, enabling agents to run in optimized environments.
1.3.3 Unification across Frameworks
  • Common Format: Agent Spec’s common format simplifies integration of agents and enables easier agent sharing across platforms.
  • Knowledge Exchange: The unified declarative format that Agent Spec provides helps developers as well as maintainers of agents to efficiently exchange solutions and ideas.
  • Conformance Test Suite: The Agent Spec conformance test suite offers a comprehensive set of tests that cover the various functionalities and components available within the specification, such as Agent, Flow, and ToolNode. It includes illustrative validation examples of Agent Spec YAML configurations and verifies that these configurations produce consistent execution outcomes and behaviors across multiple runtime environments. This approach ensures interoperability and correctness of implementations supporting the Agent Spec specification.
2 Open Agent Specification Design
This section provides an overview of Agent Spec’s declarative model. It details how Agent Spec functions across different frameworks and outlines implementation specifics, including the SDK structure, supported programming languages, agent definition using the SDK, and the serialization/deserialization of agents and workflows.
Agent Spec is intended to be a portable, platform-agnostic configuration language that allows agents and agentic systems to be described with sufficient fidelity. It defines the conceptual objects, called components, that compose agents in typical agentic systems, including the properties that determine the components’ configuration, and their respective semantics.
Runtimes implement the Agent Spec components for execution with agentic frameworks or libraries. Agent Spec would be supported by SDKs in various languages (e.g., Python), enabling serialization and deserialization of Agents to YAML, or creating them from object representations with assurance of conformance to the specification.
An outline of the components involved in the Agent Spec ecosystem is depicted in Fig. 1.
Table 2: Agent Spec aims to do for AI agents what ONNX did for ML models
Feature ONNX (ML Models) Agent Spec (AI Agents)
Scope
Unified representation of
ML models, allowing portability
across different deep learning
frameworks.
Unified representation of AI agents
and workflows, enabling them to
run across diverse agentic frameworks.
Portability
ONNX allows ML models to be
trained in one framework (e.g.,
PyTorch, TensorFlow) and
executed in another.
Agent Spec enables AI agents to be built in
one framework (e.g., LangGraph [10], AutoGen [6],
OCI Agents [4]) and run in another without modification.
Unification
Provides a common declarative
format which is versioned as
an Opset.
Provides a common agent
configuration format.
Extensibility
Supports various ML
operations and optimizations
Defines modular components for scalable agent systems.
Figure 1: Agent Spec’s Design
2.1 Overview of Agent Specification
The Open Agent Specification defines structure and behavior of the conceptual building blocks, called components, that make up agents in typical agent-based systems. This structure can be serialized into a serialization language, like JSON or YAML. An overview of the language specification of Agent Spec follows, while the full specification can be accessed at https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html.
2.2 Components
In Agent Spec, components are the building blocks which cover the commonly used elements in agentic systems, such as (but not limited to) LLMs and tools.
Describing these components with specific types enables:
  • Type safety + Usage hints for GUIs
    • Example: if a component’s property is expecting an LLM, only allow LLMs to be connected to it.
    • Example: highlight valid connections from one component to others in a graphical programming interface for agents.
  • Static analysis + validation
    • Examples: if a component is a Flow, it must have a start node. If a component is an Agent, it must have an LLM.
  • Ease of programmatic use
    • Component families each have corresponding class definitions in the Agent Spec SDK. This allows consumers (agent execution environments, editing GUIs) to utilize concrete classes, rather than having to implicitly understand which type has which properties.
2.2.1 Base Component
The basic block in Agent Spec is a Component, which by itself can be used to describe any instance of any type, guaranteeing flexibility to adapt to future changes.
Note that Agent Spec does not need to encapsulate the implementation code it describes; it simply needs to be able to express enough information to instantiate a uniquely identifiable object of a specific type with a specific set of property values.
A metadata field contains all additional information that could be useful in different usages of the Agent Specification. For example, GUIs will require including some information about the components (e.g., position coordinates, size, color, notes, …), so that they will be able to visualize them properly even after an export-import process. Metadata is optional and set to null by default.
Symbolic references (configuration components)
When a component references another component entity (for example as a property value), there is a simple symbolic syntax to accomplish this, "\$component_ref:" followed by the id of the other component.
This type of relationship is applicable to both Agent components as well as Flow components (a fixed flow component may also use an LLM or memory component).
We do not need a separate object for this type of reference, as the reference is explicitly assigned to some property or parameter.
1"\$component_ref:{COMPONENT_ID}"
Input/output schemas
Components might need some input data to perform their task and expose some output as result. These inputs and outputs must be declared and described in Agent Spec, so that users, developers and other components of the agent are aware of what is exposed as inputs and outputs respectively. Note that we do not add input/output schemas directly to the base Component class, as there are a few cases where they do not really apply (for example LLMConfig and edges, please see the next sections).
Input and output properties
Input and output schemas are used to define which values the different components accept as input or provide as output. We call these values "properties".
The term "properties" originates from JSON schema [5]. In the Agent Spec specification, we rely on this widely accepted and consolidated format. For more information about JSON schema definition, please check the official website at https://json-schema.org/understanding-json-schema.
In Agent Spec, a valid schema requires specific attributes:
  • title: the name of the property
  • type: the type of the property, following the type system described below
Additionally, users can specify:
  • default: the default value of this property
  • description: a textual description for the property
These are the minimal set of attributes (or annotations, as per JSON schema terminology) that we require to be supported for Agent Spec compatibility.
Type system
We rely on the typing system defined by the JSON schema format (see https://json-schema.org/understanding-json-schema/reference/type). At its core, JSON Schema defines the basic types available in most programming languages, though they may go by different names.
Inputs and outputs of nested components
In case of nested components (for example using an Agent component inside a Flow, or a Flow inside another Flow, or just a step in a Flow), the wrapping component is supposed to expose a (sub)set of the inputs/outputs provided by the inner components together with additional inputs/outputs it might generate as shown in Fig. 2. We require to replicate the JSON schema of each value in the inputs/outputs lists of every component that exposes it, hence, if a wrapper component exposes some inputs and outputs of some of its internal components, it will have to include them in its own list of inputs/outputs as well. This makes the specification of the components a bit more verbose, but clearer and more readable. Parsers of the specification are required to ensure the consistency of the input/output schemas defined in it.
Figure 2: Inputs and outputs of nested components, as well as those of their wrapping component
In the example above the wrapping Flow component (see sections below for more details about Flows) requires two inputs that coincide with two of the inputs exposed by the nested components and exposes one output among those exposed by the inner components in addition to one it computes itself.
Validation of specified inputs/outputs
Some input and output schemas are automatically generated by components based on their configuration. However, for the sake of clarity and readability of Agent Spec, we require that the descriptions of inputs and outputs are always reported in the Agent Specification. Consequently, when a serialized specification is read and imported by a runtime or an SDK, the inputs and outputs configuration of each component must be validated against the one generated by its configuration (if any).
Specifying inputs through placeholders
Some components might infer inputs from special parts of their attributes. For example, some nodes (e.g., the LLMNode) extract inputs from their configuration attributes, so that their values can be adapted based on previous computations We let users define placeholders in some attributes, by simply wrapping the name of the property among a double set of curly brackets. The node automatically infers the inputs by interpreting those configurations and the actual value of the input/output will be replaced at execution time. Whether an attribute accepts placeholders depends on the definition of the component itself. For example, setting the prompt_template configuration of an LLMNode to "You are a great assistant. Please help the user with this request: {{request}}" will generate an input called "request" of type string.
2.2.2 Agent
The Agent is the top-level component that holds shared resources such as conversation memory and tools. It also represents the entry point for interactions with the agentic system. Its purpose is to fill out the values for all the properties defined in the outputs attribute, with or without any interaction with the user.
It is important to have a separate definition for Agents as components, so that the same Agent component can be reused several times in a Flow or as part of a more complex agent design, without replicating its definition multiple times.
2.2.3 LLM
Large Language Models (LLM) are used in agents as generative models. To embed them into the agent, they need to be configurable flexibly with configuration parameters like the connection details, model id, the generation parameters, and more.
By default, only the generation parameters used when prompting the LLM must be specified. Specific extensions of LLMConfig for the most common model providers can be derived.
Indeed, this is not limited to locally hosted models; in principle, it also enables the configuration of API endpoints such as those provided by OpenAI, Anthropic, Mistral, and others.
2.2.4 Tool
A tool is a procedural function or a Flow that can be made available to an Agent component to execute and perform tasks. The Agent component can decide to call a tool based on its signature and description. In a fixed Flow context, a node would need to be configured to call a specific tool.
In Agent Spec, we differentiate tools according to where the actual functionality is executed or run and specify three corresponding types:
  • ServerTools are executed in the same runtime environment where the agent is being executed. The definitions of these tools therefore must be available to the agent’s environment. It is expected that this type of tool will be limited in number, and generally in functionality.
  • ClientTools are not executed by the runtime. The client environment must execute the tool and provide the results back to the runtime (like OpenAI’s function calling model).
  • RemoteTools are run in an external environment and are triggered by an RPC or REST call from the agent runtime.
Agent Spec is not supposed to provide an implementation of how these types of tools should work but provides their representation, so that they can be correctly interpreted and ported across platforms and languages. In principle a Tool is a function that is called providing some parameters (i.e., inputs), performs some transformation, and returns the result (i.e., outputs). The function has a name, and a description that can be used by an LLM for context when it is relevant to perform a task. Therefore, we define Tool as an extension of ComponentWithIO.
Tools may specify multiple outputs. In this case, the expected return value of the tool is a dictionary where each entry corresponds to an element specified in the output. The key of the dictionary entry must match the name of the property specified in the output. The runtime should parse the dictionary to extract the different outputs and bind them correctly. While ServerTool and ClientTool do not require specific additional parameters, the
RemoteTool requires including also the details needed to invoke the remote tool call.
An important security consideration for Agent Spec is the exclusion of arbitrary code from the agent specification. The specification is limited to a comprehensive description of the tool—including its attributes, inputs, outputs, and other relevant details—but does not contain any executable code.
2.2.5 Flow
Flows (or graphs) are directed, potentially cyclic workflows. They can be thought of as "subroutines" that encapsulate consistently repeatable processes. As such, they provide more determinism and reliability compared to puristic Agent components at the cost of flexibility.
Flows are executable as tools by Agent components. Each Flow requires 0 or more inputs (any input edges expressed on the starting node of the graph) and may produce 0 or more outputs (any output edges expressed by the terminal node of the graph - note that a graph can have more than one terminal node, in the case of branching logic). As such, the I/O surface of a Flow is directly mutually exchangeable with that of a tool.
Flow objects specify the the directed graph’s entry point, nodes, and edges. Components inherent in the Flow are described in subsequent sections.
A Flow exposes the set of outputs that are available in all EndNode objects, and all inputs that are defined in its StartNode.
Relationships / Edges
For Flows, Agent Spec defines separate relationships (edges) for both properties and control flow. This allows Agent Spec to unambiguously handle patterns such as conditional branches and loops, which are challenging to support with data-only flow control.
Relations are only applicable for Flows. All relations express a transition from some source node to some target node.
2.2.6 ControlFlowEdge
A control relationship defines a potential transition from one branch of a node to another. The actual transition to be taken during a given execution is defined by the implementation of a specific node.
The from_branch attribute of a ControlFlowEdge is set to null by default, which means that it will connect the default branch (i.e., “next” ) of the source node.
2.2.7 DataFlowEdge
Considering an I/O system, a component id alone is not sufficient to identify a data relationship. To this purpose, a data relationship is needed to define what output property value should be used to fill an input property.
A component may take multiple input parameters and/or produce multiple output values. We must be able to determine what value or reference is mapped to each input and where each output is used. Destinations (Inputs of the destination node) can be connected to outputs of another node or to static values. Sources (Outputs of the source node) can be connected to inputs of another node or left unconnected.
Connecting multiple data edges
(a) Control flow with two branches that converge into the same node with a property filled based on the branch taken
(b) Control flow with a self-loop and a property updated by consequent execution of the same node
Figure 3: Possible connections formed by ControlFlowEdges and DataFlowEdges. ControlFlowEdges are shown as solid lines, DataFlowEdges with dotted lines.
In the definition of a Flow, it is allowed to connect two or more different data outputs to the same input as shown in Fig. 3. The most recently executed node with an output connected to the input takes precedence. In other words, the behavior is like having the input exposed as a public variable and every node that has an output connected to that input updates its value.
For the control flow, instead, edges just define which are the allowed "next step" transitions: multiple connections from the same outgoing control flow branch are not allowed, and they do not involve any ambiguity (note that we do not enable parallelism through edges).
Optionality of data relationships
Some frameworks do not require specifying the data flow explicitly as they assume that all the properties are publicly exposed in a shared "variable space" that any component can access, and the read/write strategy is simply name based. In practice, this means that:
  • When a component has an input with a specific name, it will access the public variable space to read the value of the variable with that name.
  • When a component has an output with a specific name, it will write (or overwrite if it already exists) into the public variable space the variable with that name.
We let users adopt this type of I/O system by setting the data_flow_connections parameter of the Flow to None. In this case, the name-based approach explained above will be adopted, by using the shared variable space described above as the public space where values are published.
Note that this name-based approach can be expressed by explicitly defining data-flow connections (but the opposite is not possible) by connecting all the outputs to the inputs with the same name, following the control flow connections in the Flow to account for overwrites.
Flow state
We define as "flow state" the set of all values that describe the current stage of execution within the flow (the point of execution reached). Users have the ability to specify which set of values within the flow state are publicly exposed to all nodes.
These values are treated differently compared to values that are exposed through inputs and outputs of nodes:
  • They are always available in every node and do not need to be connected by a DataFlowEdge.
  • If a component writes an output which name matches a flow state value, the respective value in the flow state is overwritten.
Conversation
At the core of the execution of a conversational agent, there is the conversation itself. The conversation is implicitly passed to all the components throughout the flow. It contains the messages being produced by the different nodes: each node in a flow and each agent can append messages to the conversation if needed. Messages have, at least, a type (e.g., system, agent, user), content, the sender, and the recipient.
2.2.8 Node
A Node is a vertex in a Flow and is derived from ComponentWithIO. Agent Spec provides a library of nodes designed to streamline workflow orchestration, enabling rapid development and seamless integration for a wide range of applications:
  • LLMNode: uses a LLM to generate some text from a prompt.
  • APINode: performs an API call with the given configuration and inputs.
  • AgentNode: runs a multi-round conversation with an Agent component, used to better structure agents and easily reuse Agent components.
  • FlowNode: runs a Flow, used to better structure agents and easily reuse Flows.
  • MapNode: performs a map-reduce operation on a given input collection, applying a specified Flow to each element in the collection.
  • StartNode: entry point of a Flow.
  • EndNode: exit point of a Flow.
  • BranchingNode: allows conditional branching based on the value of an input.
  • ToolNode: executes a tool.
A more detailed description of each node type is given in https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html#standard-library-of-nodes-to-use-in-flows.
2.3 Open Agent Specification Model
The objective of Agent Spec is to be agnostic with respect to any agentic framework and the programming language used to build and execute an agent. For this reason, an agent represented with Agent Spec should be exportable to, and importable from a representation that is specified in a common serialization language. To this end, Agent Spec adopts JSON as the designated language for the serialization of its components.
All building blocks of Agent Spec are designed with the goal of being trivially serializable. It follows that the Agent Specification of a component can be obtained by serializing its definition according to the JSON specification. In particular, the serialization of a component includes its attributes, plus an additional field called component_type that will be used at deserialization time to infer the correct component type to deserialize.
Note that, as defined in the Agent Spec language, components can be declared only once in the representation and then referenced according to the rules described in the Symbolic references section of this document.
A few examples of component JSON serializations are given at https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html#language-examples.
2.4 How Agent Spec Works Across Frameworks
Agent Spec serves as an abstraction layer above framework-specific implementations, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints.
Hence, the goal of Agent Spec is to define behavioral patterns for the different components that should be implemented according to the specificities of the different agentic frameworks. An Agent Spec runtime for a specific framework should be able to read the Agent Specification of a component (e.g., a Flow or an Agent), and then execute its logic, adhering to the behavior described in the Agent Spec.
As an alternative, if the agentic framework - used to implement the runtime - offers a serialization mechanism, it is also possible to build a translation script between the specification of an agent and the equivalent framework specific representation, then execute the runtime on the latter.
2.5 Implementation Details
2.5.1 Agent Spec SDKs
To facilitate the process of building programmatically framework-agnostic agents, Agent Spec SDK packages can be conveniently implemented. These SDKs should provide two functionalities:
  • Building Agent Spec component abstractions through the implementation of the relevant interfaces, in conformity with the Agent Spec.
  • Import and export these abstractions from and to their respective serialized version in JSON format.
As part of the Agent Spec effort, we provide Agent Spec’s Python SDK that we call PyAgentSpec. PyAgentSpec allows building Agent Spec-compliant agents in python. Users can create their own assistants using component classes that replicate the interface and behavior defined by Agent Spec, which can finally be exported in JSON format. Here is an example of how the definition of a simple agentic Flow prompting an LLM with a custom user input and returning the LLM’s output would look like in PyAgentSpec.
1from pyagentspec.property import Property
2from pyagentspec.flows.flow import Flow
3from pyagentspec.flows.edges import ControlFlowEdge, DataFlowEdge
4from pyagentspec.flows.nodes import LlmNode, StartNode, EndNode
5from pyagentspec.llms import VllmConfig
6llm_config = VllmConfig(
7 name="<Your Model Name Here>",
8 url="<url.of.your.llm.deployment:port>",
9 model_id="<provider/model-identifier>",
10)
11prompt_property = Property(
12 json_schema={"title": "prompt", "type": "string"}
13)
14llm_output_property = Property(
15 json_schema={"title": "llm_output", "type": "string"}
16)
17start_node = StartNode(name="start", inputs=[prompt_property])
18end_node = EndNode(name="end", outputs=[llm_output_property])
19llm_node = LlmNode(
20 name="simple llm node",
21 llm_config=llm_config,
22 prompt_template="{{prompt}}",
23 inputs=[prompt_property],
24 outputs=[llm_output_property],
25)
26flow = Flow(
27 name="Simple prompting flow",
28 start_node=start_node,
29 nodes=[start_node, llm_node, end_node],
30 control_flow_connections=[
31 ControlFlowEdge(name="start_to_llm", from_node=start_node, to_node=llm_node),
32 ControlFlowEdge(name="llm_to_end", from_node=llm_node, to_node=end_node),
33 ],
34 data_flow_connections=[
35 DataFlowEdge(
36 name="prompt_edge",
37 source_node=start_node,
38 source_output="prompt",
39 destination_node=llm_node,
40 destination_input="prompt",
41 ),
42 DataFlowEdge(
43 name="llm_output_edge",
44 source_node=llm_node,
45 source_output="llm_output",
46 destination_node=end_node,
47 destination_input="llm_output"
48 ),
49 ],
50)
To export the Flow that was just built in JSON format, the code is the following.
1serializer = AgentSpecSerializer()
2serialized_flow = serializer.to_json(flow)
The full documentation of PyAgentSpec, including installation instructions, APIs documentation and examples, is available at https://oracle.github.io/agent-spec/index.html.
2.5.2 Agent Spec SDK Plugins
The PyAgentSpec SDK also supports the concept of plugins, because we expect that teams using Agent Spec may want to create new component types to be included in their configurations. To this end, a plugin can enable newer components in addition to the standard components of Agent Spec, simply by specifying the list of component types that the plugin includes and by providing the serialization and deserialization logic for the plugin. Such plugins can be implemented for components not yet included in the first version of Agent Spec, such as multi-agent patterns, or planning and memory components. See below the abstract interface to be implemented for a serialization plugin (the deserialization interface follows the same logic).
1from abc import ABC, abstractmethod
2from typing import Any, Dict, List
3
4from pyagentspec.component import Component
5from pyagentspec.serialization.serializationcontext import SerializationContext
6
7class ComponentSerializationPlugin(ABC):
8 """Base class for Component serialization plugins."""
9
10 @property
11 @abstractmethod
12 def plugin_name(self) -> str:
13 """Return the plugin name."""
14 pass
15
16 @property
17 @abstractmethod
18 def plugin_version(self) -> str:
19 """Return the plugin version."""
20 pass
21
22 @abstractmethod
23 def supported_component_types(self) -> List[str]:
24 """Indicate what component types the plugin supports."""
25 pass
26
27 @abstractmethod
28 def serialize(
29 self, component: Component, serialization_context: SerializationContext
30 ) -> Dict[str, Any]:
31 """Serialize a component that the plugin should support."""
32 pass
2.5.3 Agent Spec Runtime Adapters
A runtime implementation makes Agent Spec components abstractions runnable by implementing the behavioral logic described in the Agent Spec language.
As previously discussed, there are several ways to implement runtimes for Agent Spec. Ideally, a runtime implementation is an agentic framework that natively supports Agent Spec: it should read an Agent Spec JSON specification and transform it directly into an equivalent component that can be executed.
As an example of an Agent Spec runtime implementation, we provide a package called WayFlow, which is a powerful, intuitive Python library for building sophisticated AI-powered assistants. It includes a library of plan steps to streamline the creation of AI-powered assistants, supports reusability, and it is ideal for rapid development. The WayFlow framework provides a serialization and deserialization interface to import and export Agent components according to the Agent Spec. It is a reference framework for Agent Spec and our runtime of reference, with native support for all Agent Spec Agents and Flows.
When an agentic framework does not allow reading Agent Specifications directly, a translation layer that transforms the Agent Specifications into a framework-specific representation could be adopted. This layer, which we can call Agent Spec Runtime adapter, loads an Agent Specification and transforms it into an equivalent component of the specific agentic framework.
Runtime adapters enable existing agentic frameworks that do not support Agent Spec natively executing Agent Specifications. For example, we also provide runtime adapters for some of the most popular agentic frameworks, including, but not limited to, LangGraph and Microsoft AutoGen. These adapters load Agent Specifications taking advantage of the Agent Spec Python SDK, and they transform the Agent Spec components into the respective versions defined in the agentic framework of interest.
2.6 Performance Evaluation & Benchmarks
Agent Spec enables key aspects of productizing agent-driven solutions, including comparison and evaluation of agent task performance across different executing frameworks and models, as well as optimizations (including costs) of execution for underlying hardware and frameworks. By covering commonly accepted concepts in agentic systems, such as workflows and agents, Agent Spec allows developers to map these concepts to various execution frameworks, facilitating comparison of framework specific task performance. Compute optimization is also possible, as Agent Spec enables hardware-optimized and framework-optimized implementation of agentic design patterns.
2.6.1 Portability Analysis
Agentic AI is rapidly evolving, and a unified benchmark for assessing agent reliability, particularly in enterprise settings, is currently lacking. Future updates for Agent Spec report will include empirical results demonstrating the reliability of Agent Spec-defined agents across different frameworks.
https://oracle.github.io/agent-spec/howtoguides/howto_execute_agentspec_across_frameworks.html provides an example of how an agent defined with Agent Spec can be run in different frameworks. In this Retrieval Augmented Generation case study, an agent utilizing the same tools is instantiated in AutoGen, LangGraph, and WayFlow. In a practical setting, a specific mapping from Agent Spec to each framework would be implemented.
2.6.2 Use Cases & Real-World Applications
A common use case of Agent Spec is the porting of agent-based assistants from one framework to another. Let’s take as an example a team that implemented agent-based workflows using AutoGen. As business and process needs to change, evaluations are conducted to consider alternative frameworks. Due to AutoGen’s specific architecture, migrating to another agentic framework proves to be complex with significant rewriting overhead. In this example, Agent Spec enables the definition of business unit-specific agentic systems using a declarative representation. This allows for conversion to specific framework formats or flexible execution across different agentic frameworks with appropriate adapters, enabling straightforward comparison of task performance and runtime cost depending on the execution framework. Consequently, utilizing Agent Spec throughout the agent design and deployment process reduces complexity, costs, implementation effort, and potential inconsistencies.
Another use case involves developing and deploying Agents across varied compute environments, such as cloud platforms and on-premise datacenters. In the cloud, Agents may run virtualized or serverless as part of a scalable system design. In an on-premise scenario, they may run within an application server. Agent Spec’s framework and runtime agnosticism enables development of the Agent in a local development environment, followed by flexible deployment across either of these scenarios.
A different compelling real-world application involves advanced multi-agent systems in the domain of financial crime compliance. By integrating generative AI agents, it enables to automate key investigative processes - supplementing human investigators with narrative-driven analyses, uncovering critical insights, and recommending evidence-based actions. In this context, a comprehensive Anti-Money Laundering platform may comprise specialized agents tasked with sanctions screening, transaction monitoring for suspicious activity, and the generation of enhanced due diligence reports. These sub-agents are often developed by independent teams using a variety of frameworks and technologies, and may need to be deployed across heterogeneous environments, including cloud, on-premises, or even within database systems. Leveraging Agent Spec’s framework-agnostic and declarative approach, development teams can seamlessly integrate and orchestrate these diverse agent workflows, reducing integration overhead and enabling consistent, scalable deployment. This holistic and automated strategy minimizes manual effort, decreases operational risk, and empowers financial institutions to respond more rapidly and consistently to evolving financial crime patterns, regardless of the underlying deployment environment.
2.7 Positioning in the Ecosystem
Agent Spec aims to streamline the architecture and design of agentic assistants and workflows, serving as a unified representation that abstracts away implementation details of specific agentic frameworks. It is a portable configuration language that describes agentic design patterns, components and expected behavior. However, Agent Spec is not the only effort aimed at unifying the different parts that compose a common agentic ecosystem.
  • The Model Context Protocol (MCP) [7], introduced by Anthropic, standardizes resource and data provision to agents. Resources are made available with a client/server based RESTful API.
  • Google’s Agent2Agent Protocol [2] and BeeAi’s Agent Communication Protocol (ACP) [3] propose standardized APIs for distributed agent communication.
  • Nvidia recently announced NeMo Agent toolkit which is a framework-independent library that treats “every agent, tool, and agentic workflow as a function call—enabling composability between these agents, tools, and workflows”[1]. The idea is to compose existing agentic solutions and resources, potentially written using different frameworks, into new systems.
While protocols like MCP and A2A standardize tool or resource provisioning as well as inter-agent communication, Agent Spec complements these efforts by enabling standardized configuration of components related to agentic system design and execution in general, as shown in Fig. 4. Hence, Agent Spec provides strong synergies with other standardization efforts through its abstract definition of agentic system design. In the context of ongoing unification efforts within the agentic system landscape, Agent Spec aims to serve as the "common foundation" that connects these initiatives, enhancing their effectiveness and fostering a more cohesive ecosystem. Support for the aforementioned protocols may be added in future versions of Agent Spec as new components.
Figure 4: Agent Spec complements other standardizations, such as MCP or A2A
2.8 Future Roadmap & Enhancements
2.8.1 Upcoming Features
Expanding the Open Agent Specification language We are looking forward to extending Agent Spec with new concepts and components in the upcoming versions of Agent Spec. A few notable ones include:
  • Memory
  • Planning
  • Datastores
  • Support for remote agents (e.g., A2A agents)
  • Support for new types of tools (e.g., MCP tools)
Extending runtimes support
We want to foster the implementation of Agent Spec runtime adapters for the popular agentic frameworks in the vibrant AI community. At the same time, we aim to offer a conformance test suite for runtimes to ensure consistent behavior in accordance with the Agent Specification.
Improving user experience
Even though Agent Spec SDKs simplify the creation and validation of Agent Specifications, creating agents by writing code might be complex and error-prone. We are working on a Drag&Drop UI that will allow users to build their own agents in a visual interface and then export the Agent Specification to be used in the framework of their choice via a runtime adapter.
2.8.2 Community & Ecosystem Growth
We strongly believe Agent Spec will contribute to streamlining agent development and open-sourcing its specification to encourage community contributions. This will foster a rich ecosystem of interchangeable and reusable designs and tools for agentic AI development. Integration of Agent Spec by other third-party open-source projects as an agent design language and specification will benefit the open-source community, which is currently evolving and lacks established common unified representation. In summary, the key aspects of impact on community growth are:
  • Design patterns that can be easily shared & reused to collaboratively build increasingly complex solutions.
  • Enable open exchange of building blocks and agentic patterns.
  • Encourage contributions by the open-source community.
To help guide the evolution of the Agent Specification in alignment with its core design principles, we will establish a steering committee. Similar to ONNX principles, this committee will work closely with the community to review proposed contributions, foster open discussion, and publish versioned updates.
3 Contributors and Acknowledgements
Please cite this work as “Open Agent Specification (2025)”.
Yassine Benajiba, Cesare Bernardis, Vladislav Blinov, Paul Cayet, Hassan Chafi, Abderrahim Fathan, Louis Faucon, Damien Hilloulin, Sungpack Hong, Ingo Kossyk, Rhicheek Patra, Sujith Ravi, Jonas Schweizer, Jyotika Singh, Shailender Singh, Xuelin Situ, Weiyi Sun, Jerry Xu, Ying Xu. [^1] Author list is sorted alphabetically.
We are also thankful to our Oracle colleagues who engaged in active discussions during the design of Agent Spec and provided valuable feedback that helped us refine its specifications.

### v3

@@
 This technical report provides an overview of the technical foundations of Agent Spec, including motivation, benefits, and future developments.
+Agent Spec also introduces a standardized Evaluation harness to assess agent behavior and agentic workflows across runtimes - analogous to how HELM and related harnesses standardized LLM evaluation - so that performance, robustness, and efficiency can be compared consistently across frameworks.
+We demonstrate this using four distinct runtimes (LangGraph, CrewAI, AutoGen, and WayFlow) evaluated over three different benchmarks (SimpleQA Verified, $\tau^{2}$-Bench and BIRD-SQL).
 1 Introduction
@@
 Agent Spec serves as an abstraction layer above framework-specific specifications, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints.
+It also serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across different frameworks (e.g., LangGraph, CrewAI, AutoGen, and WayFlow).
+Using this capability, we report initial results on representative tasks from three benchmark datasets (SimpleQA Verified [4], $\tau^{2}$-Bench [2] and BIRD-SQL [14]).
 1.1 Background & Motivation
@@
 Similar to ONNX principles, this committee will work closely with the community to review proposed contributions, foster open discussion, and publish versioned updates.
+3 Evaluation & Benchmarking
+Agent Spec supports key aspects of the productization of agent-driven solutions, including the comparison and evaluation of agent task performance across different execution frameworks and model configurations, as well as optimization of execution costs and efficiency across different environments.
+By encompassing widely accepted concepts in agentic systems, such as workflows and agents, Agent Spec provides a unifying abstraction that enables developers to map these concepts consistently across diverse execution frameworks.
+This facilitates direct performance benchmarking and comparative analysis of framework-specific task execution.
+Furthermore, compute optimization is also possible, as Agent Spec enables hardware-optimized and framework-optimized implementation of common agentic design patterns.
+Please see Appendix C for additional details on real-world use cases & business applications.
+We note that the goal of this evaluation is not to outperform existing implementations of agents, but to show that Agent Spec enables cross-framework portability and reusability of non-trivial, competitive agents.
+3.1 Evaluation methodology
+We evaluate Agent Spec on three popular benchmarks: SimpleQA Verified [4], $\tau^{2}$-Bench [2] and BIRD-SQL [14].
+For each benchmark, we select one or more Agent Spec JSON configurations and execute the same specification across four framework runtimes: AutoGen, CrewAI, LangGraph, and WayFlow.
+This approach allows direct comparison of behavioral consistency, correctness, and latency across heterogeneous frameworks.
+In this evaluation, the same agent utilizing identical tools is instantiated in AutoGen, CrewAI, LangGraph, and WayFlow.
+We report F1 scores for SimpleQA Verified (exact-match over verified answers), EX% for BIRD-SQL (execution accuracy for generated SQL compared to ground-truth SQL), and Passˆk for $\tau^{2}$-Bench (probability to get all k attempts correct out of k).
+Query time is reported as mean ± standard deviation over all records.
+SimpleQA Verified.
+We evaluate on the full set of 1,000 questions using two representative assistants: (i) a ReAct-style agent equipped with a web-browsing tool, and (ii) a flow-based agentic RAG solution that applies question decomposition, web-browsing, and self-reflection to answer complex user queries [^1] Note that the browse tool makes the task [4] easier, but it is acceptable since we are demonstrating Agent Spec’s functional effectiveness instead of competing on the benchmark leaderboard..
+We also include the performance of a single LLM call as a baseline.
+Bird-SQL
+We evaluate execution accuracy on the 1,534 questions from BIRD-SQL dev with two types of assistants.
+The first assistant is a ReAct-style agent equipped with a tool to validate that a query can be executed, and a "thinking" tool to give opportunities to the agent to think between tool calls.
+The second assistant is a flow-based assistant composed of a planning step, a step to generate the SQL query and a final post-processing step to self-reflect on the generated query.
+We also include the performance of a single LLM call as a baseline.
+$\tau^{2}$-Bench.
+A simulation environment designed to evaluate conversational customer service agents in a dual-control setting, where both the agent and a user simulator interact within realistic domains such as airline, retail, and telecom.
+For each domain the benchmark provides a set of tools to configure the agent, and requests are given to the agent via a user simulation.
+Both the actual actions of the agent and the information communicated to the user are scored by the benchmark.
+We assess performance in the retail domain using a max-steps limit of 40 and GPT-4.1 as the user model.
+For the ReAct-style agent, we evaluate both GPT-4.1 and GPT-4.1 mini as the underlying models.
+3.2 Results and Insights
+Agent Spec successfully instantiates functional agents across all tested frameworks, confirming consistent behavior and validating its framework-agnostic portability & reusability.
+This confirms that a single declarative specification can reproduce agent execution across heterogeneous runtimes.
+SimpleQA Verified.
+On SimpleQA (Table 3), the large gap between simple prompting of GPT-4.1 mini and the ReAct agent implementation (both based on the same model) shows that all frameworks gain substantially from the use of web search tools.
+Even a compact model like GPT-4.1 mini can successfully orchestrate ReAct agents.
+Using a more powerful model to orchestrate the agent does not yield meaningful quality improvements.
+Latency differences mainly arise because the Llama model runs locally while GPT-4.1 mini depends on a remote API with variable response times.
+The flow-based solution leads to higher accuracy thanks to decomposition and self-reflection.
+High variance in time per query is connected to the different number of reasoning iterations (see Appendix D.1 for more information on the flow implementation).
+Differences in ReAct agent performance reveal interesting insights into each framework’s implementation, which in turn affect the final outcomes.
+  • WayFlow and LangGraph show similar behaviors, with WayFlow being slightly more accurate but slower than LangGraph.
+  • AutoGen, a multi-agent system–oriented framework, performs less effectively in this single-agent scenario.
+Its lack of support for non-native tool calling further limits its performance compared with other frameworks using the Llama model.
+  • CrewAI achieves the highest F1 score but at the cost of significantly higher response times between 55% (with GPT-4.1 mini) to 140% slower (with Llama 3.3 70B) than competitors and higher setup complexity, as it requires additional directives to create the agent’s task.
+Table 3: Benchmark results on SimpleQA Verified. Flow-based results for AutoGen and CrewAI are not reported since they do not support non-agentic flow execution.
+Agent Setup Runtime F1-score (%) Time/query (s)
+GPT-4.1 mini - 17.8 $1.06\pm 0.41$
+ReAct Agent (GPT-4.1 mini) WayFlow 79.0 $6.22\pm 5.24$
+LangGraph 75.5 $4.57\pm 6.21$
+AutoGen 67.9 $4.64\pm 1.64$
+CrewAI 81.1 $7.21\pm 6.74$
+ReAct Agent (Llama 3.3 70B) WayFlow 79.2 $4.60\pm 0.92$
+LangGraph 77.6 $4.10\pm 0.91$
+AutoGen 54.9 $3.38\pm 0.31$
+CrewAI 82.3 $8.12\pm 3.84$
+Flow-based agentic RAG (GPT-4.1 mini) WayFlow 88.1 $23.19\pm 43.25$
+LangGraph 85.3 $23.21\pm 55.53$
+BIRD-SQL.
+On BIRD-SQL (Table 4), the performance gap between the different frameworks is more narrow compared to SimpleQA and $\tau^{2}$-Bench, which is to be expected given that NL2SQL tasks require short trajectories to be solved.
+CrewAI is still significantly slower compared to the three other runtimes.
+Giving tools to the Agent does not improve over the baseline performance but the Plan-Generate-Reflect flow consistently outperforms the baseline.
+Table 4: Benchmark results on BIRD-SQL. Flow-based results for AutoGen and CrewAI are not reported since they do not support non-agentic flow execution.
+Agent Setup Runtime EX (%) Time/query (s)
+GPT-4.1 mini - 50.1 $4.26\pm 2.25$
+GPT-4.1 - 54.4 $2.93\pm 4.68$
+ReAct Agent (GPT-4.1) WayFlow 54.1 $7.16\pm 2.72$
+LangGraph 53.7 $5.68\pm 3.64$
+AutoGen 54.1 $6.06\pm 2.97$
+CrewAI 54.3 $12.20\pm 6.62$
+NL2SQL Flow (GPT-4.1) WayFlow 55.5 $9.45\pm 5.56$
+LangGraph 55.9 $7.54\pm 4.11$
+$\tau^{2}$-Bench.
+Performance differences on $\tau^{2}$-Bench (Table 5) stem from variations in prompt-handling strategies across frameworks which wrap the system prompt and add additional instructions in distinct ways.
+These frameworks also differ in how they request tool usage from the hosted model.
+Such differences are influenced by how each model is hosted and how it manages API calls and responses.
+Notably, Agent Spec exposes these variations, enabling consistent cross-framework evaluation.
+Additionally, we observe the lowest scores for the CrewAI runtime, which can be attributed to its design focus on multi-agent coordination for single-task problem-solving, rather than on conversational interactions with users as evaluated by $\tau^{2}$-Bench.
+Table 5: Benchmark results on $\tau^{2}$-Bench. Experiments are grouped by the base LLM used.
+Agent Setup Runtime Passˆ1 / ˆ4 (%) Time/query (s)
+ReAct Agent (GPT-4.1) WayFlow 64.7 / 43.0 $43.0\pm 13.8$
+LangGraph 71.5 / 51.8 $37.5\pm 12.4$
+AutoGen 73.5 / 52.6 $55.5\pm 17.8$
+CrewAI 62.5 / 38.6 $62.7\pm 25.4$
+ReAct Agent (GPT-4.1 mini) WayFlow 60.1 / 32.5 $40.9\pm 12.7$
+LangGraph 60.7 / 32.5 $38.4\pm 17.4$
+AutoGen 58.3 / 36.0 $47.4\pm 15.0$
+CrewAI 51.5 / 22.8 $56.1\pm 25.5$
 4 Contributors and Acknowledgements
 Please cite this work as “Open Agent Specification (2025)”.
-Yassine Benajiba, Cesare Bernardis, Vladislav Blinov, Paul Cayet, Hassan Chafi, Abderrahim Fathan, Louis Faucon, Damien Hilloulin, Sungpack Hong, Ingo Kossyk, Rhicheek Patra, Sujith Ravi, Jonas Schweizer, Jyotika Singh, Shailender Singh, Xuelin Situ, Weiyi Sun, Jerry Xu, Ying Xu.
+Yassine Benajiba, Cesare Bernardis, Vladislav Blinov, Paul Cayet, Hassan Chafi, Abderrahim Fathan, Louis Faucon, Damien Hilloulin, Sungpack Hong, Ingo Kossyk, Rhicheek Patra, Sujith Ravi, Jonas Schweizer, Jyotika Singh, Shailender Singh, Xuelin Situ, Weiyi Sun, Kartik Talamadupula, Jerry Xu, Ying Xu.
 [^2] Author list is sorted alphabetically.

### v4

@@
-Open Agent Specification (Agent Spec) Technical Report
+
 Abstract
-Open Agent Specification (Agent Spec) is a declarative language that allows AI agents and their workflows to be defined in a way that is compatible across different AI frameworks, promoting portability and interoperability within AI Agent frameworks.
-Agent Spec aims to resolve the challenges of fragmented agent development by providing a common unified specification that allows AI agents to be designed once and deployed across various frameworks, improving interoperability and reusability, and reducing redundant development efforts.
-Additionally, Agent Spec facilitates development tools and portability, allowing AI agents to be defined independently of their execution environment and enabling teams to exchange solutions without implementation-specific limitations.
-Agent Spec benefits four key groups: (i) Agent developers, who gain access to a superset of reusable components and design patterns, enabling them to leverage a broader range of functionalities; (ii) Agent framework and tool developers, who can use Agent Spec as an interchange format and therefore benefit from the support of other frameworks as well as other tools; (iii) Researchers, who can achieve reproducible results and comparability, facilitating more reliable and consistent outcomes; (iv) Enterprises, which benefit from faster prototype-to-deployment, increased productivity, as well as greater scalability and maintainability for their AI agent solutions.
-This technical report provides an overview of the technical foundations of Agent Spec, including motivation, benefits, and future developments.
+The proliferation of agent frameworks has led to fragmentation in how agents are defined, executed, and evaluated.
+Existing systems differ in their abstractions, data flow semantics, and tool integrations, making it difficult to share or reproduce workflows.
+We introduce Open Agent Specification (Agent Spec), a declarative language that defines AI agents and agentic workflows in a way that is compatible across frameworks, promoting reusability, portability and interoperability of AI agents.
+Agent Spec defines a common set of components, control and data flow semantics, and schemas that allow an agent to be defined once and executed across different runtimes.
 Agent Spec also introduces a standardized Evaluation harness to assess agent behavior and agentic workflows across runtimes - analogous to how HELM and related harnesses standardized LLM evaluation - so that performance, robustness, and efficiency can be compared consistently across frameworks.
 We demonstrate this using four distinct runtimes (LangGraph, CrewAI, AutoGen, and WayFlow) evaluated over three different benchmarks (SimpleQA Verified, $\tau^{2}$-Bench and BIRD-SQL).
+We provide accompanying toolsets: a Python SDK (PyAgentSpec), a reference runtime (WayFlow), and adapters for popular frameworks (e.g., LangGraph, AutoGen, CrewAI).
+Agent Spec bridges the gap between model-centric and agent-centric standardization & evaluation, laying the groundwork for reliable, reusable, and portable agentic systems.
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
 1 Introduction
-This technical report introduces Agent Spec, a declarative, framework-agnostic configuration language designed to define AI agents and their workflows with high fidelity.
-By providing a unified representation, Agent Spec enables seamless portability and interoperability of agents across diverse frameworks, ensuring consistent behavior and execution across various AI Agent frameworks.
-Agent Spec serves as an abstraction layer above framework-specific specifications, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints.
-It also serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across different frameworks (e.g., LangGraph, CrewAI, AutoGen, and WayFlow).
-Using this capability, we report initial results on representative tasks from three benchmark datasets (SimpleQA Verified [4], $\tau^{2}$-Bench [2] and BIRD-SQL [14]).
-1.1 Background & Motivation
-Existing agentic frameworks have their own strengths and limitations.
+The rapid emergence of AI agent frameworks has accelerated the development of systems capable of planning, invoking tools, and coordinating multi-step workflows [SAPKOTA2026103599; orclblog; xie2025ai; wolflein-etal-2025-llm].
+However, this progress has also introduced fragmentation: each framework defines its own abstractions, configuration formats, and execution assumptions [derouiche2025agenticaiframeworksarchitectures; kapoor2025ai].
+As a result, organizations struggle when porting agents across stacks, evaluating behaviors consistently, or reusing designs developed by other teams.
+These challenges hinder reproducibility, slow down prototyping-to-deployment, and complicate governance in enterprise environments [10.1145/3711896.3736570; chen2025aiopslab].
+Concurrently, several standardization efforts are shaping the agent ecosystem.
+Protocols such as Anthropic’s Model Context Protocol (MCP) [mcp] focus on resource and tool provisioning; inter-agent communication initiatives like Google’s Agent2Agent [a2a] and BeeAI’s Agent Communication Protocol (ACP) [acp] standardize message exchange.
+In addition, the Agntcy initiative [agntcy] aims to promote open interoperability and governance standards for the emerging agent ecosystem, providing shared principles for agent definition, discovery, and interaction across frameworks.
+While these efforts advance interoperability at the resource and communication layers, there still exists a critical gap at the level of agent behavior and execution semantics.
+In practice, this means there is no common standard for how control and data flows are managed, how agent state evolves, or how inputs and outputs are validated, leading to inconsistencies in agent execution behavior across different runtime environments.
+To address these gaps, we introduce Agent Spec, a declarative, framework-agnostic configuration language for defining AI agents and their workflows with high fidelity.
+By providing a unified representation, Agent Spec enables seamless reusability, portability and interoperability across diverse frameworks, ensuring consistent behavior and reproducibility.
+In practice, Agent Spec serves as an abstraction layer above framework-specific implementations, providing a unifying interface that captures agent functionality beyond individual framework constraints.
+It makes evaluation a first-class concern by defining common execution semantics and measurement protocols, so that agent and workflow performance can be compared across frameworks - similar in spirit to how HELM [liang2023holistic] and related harnesses standardized LLM evaluation.
+Like HELM, EleutherAI’s lm-evaluation-harness [eleutherai2024lmEvalHarness] and MT-Bench [bai2024mtbench101], Agent Spec centers evaluation as a core principle.
+However, while these efforts focus on model-centric standardized assessment, Agent Spec standardizes agent- and workflow-centric standardized assessment across frameworks.
+By fixing a common intermediate representation, execution semantics, and measurement protocol, Agent Spec supports reproducible, cross-platform comparisons of agent quality, safety, robustness, and efficiency, thus bridging the gap between agent portability, reusability and rigorous evaluation.
+Our contributions in this work are as follows:
+  • We present Agent Spec, a declarative, framework-agnostic language for specifying agents and workflows with high fidelity, enabling define-once, run-anywhere portability and seamless reusability across diverse runtimes.
+  • We formalize a core set of components and execution semantics - Agents, Flows, Nodes, Tools, and control/data flow edges and discuss design principles for reliable, composable, and reusable agentic systems.
+  • We release accompanying toolsets: a Python SDK (PyAgentSpec  [^1] https://pypi.org/project/pyagentspec/) for programmatic authoring and serialization, a reference runtime (WayFlow  [^2] https://github.com/oracle/wayflow) with native Agent Spec support, and runtime adapters for popular frameworks to demonstrate cross-framework execution.
+  • Agent Spec serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across different frameworks.
+Using this capability, we report initial results on representative tasks from three benchmark datasets.
+Interestingly, we observe that both LangGraph and WayFlow deliver reliable accuracy and query time stability across datasets, AutoGen outperforms on complex reasoning tasks (e.g., on $\tau^{2}$-Bench [barres2025tau2]) but with increased latency, and CrewAI’s multi-agent coordination yields limited accuracy gains despite longer query times.
+2 Background & Motivation
+2.1 Related Work
+AI agent frameworks provide the building blocks for constructing agentic systems that can reason, make decisions, and complete multi-step tasks.
+In recent years, with the advent of large language models (LLMs), several general-purpose frameworks have emerged with distinct abstractions, focus areas, and strengths, including LangChain [LangChain], LangGraph [LangGraph], GPT-Functions [openaiFunctionCalling], AutoGen [autogen], OCI Agents [ociagents], Semantic Kernel [semanticKernel], and CrewAI [crewAI].
+These frameworks can augment LLMs (e.g., through Agents, Flows) with structured planning [wang2022self], memory [wang2023augmenting], tool use [schick2023toolformer], and the ability to perform search and access external environments [zhu2023large; kim2024rada].
+LangChain [LangChain] offers a flexible, modular SDK for composing applications with LLMs, tools, long-term memory, and retrieval, and it underpins a large ecosystem of integrations.
+LangGraph [LangGraph] builds on LangChain to model agents and workflows as stateful, directed graphs with explicit control flow, enabling deterministic branching, loops, and multi-actor coordination; its associated tooling (e.g., LangGraph Studio) provides visualization and debugging support.
+AutoGen [autogen] focuses on multi-agent conversations with tool use and human-in-the-loop capabilities, providing abstractions for autonomous agent roles, inter-agent messaging, and extensible tool execution, which makes it suitable for complex, iterative workflows.
+It is designed with an open and flexible conversational communication style between agents, and its event-driven and distributed architecture makes it suitable for workflows that require long-running agents that collaborate across information boundaries with variable degrees of human involvement.
+Semantic Kernel [semanticKernel] is a model-agnostic SDK (with support across multiple programming languages) that provides planners, memories, connectors, and orchestration utilities to integrate LLMs and data sources into production applications.
+OCI AI Agent Platform [ociagents] targets enterprise-grade deployments with built-in tools and patterns (e.g., natural language SQL generation, retrieval-augmented generation), emphasizing scalability, manageability, and integration with enterprise services.
+CrewAI [crewAI] centers on multi-agent collaboration with explicit roles, tasks, and shared state, offering opinionated primitives for coordination and memory.
+Separately, GPT-Functions (tool/function calling) [openaiFunctionCalling] specifies how models such as GPT-4o can invoke user-defined tools through structured function signatures, enabling safer and more controllable LLM-tool interaction across many runtimes.
+OctoTools [lu2025octotools] proposes an extensible framework oriented toward complex reasoning with “tool cards” that standardize wrappers around heterogeneous tools to improve transparency, reduce integration errors, and simplify maintenance.
+While these frameworks advance agentic capabilities in complementary directions - ranging from SDKs to orchestration, coordination, and tool interoperability - the ecosystem remains fragmented.
+Differences in configuration models, component semantics, and execution assumptions hinder reusability, reliability, and reproducibility.
+This fragmentation motivates a unifying, declarative, and framework-agnostic specification such as Agent Spec, which provides a reusable & portable representation of agents and workflows executable across diverse runtimes.
+2.2 Motivation
+Existing agentic frameworks have their own strengths and limitations (e.g., limited tool types, lack of generalization across diverse tasks) [raheem2025agentic; huang2025ai; joshi2025review].
 While flexibility is crucial to accommodating diverse agentic design patterns, currently available frameworks do not fully support all of them.
-Table 1 outlines key objectives a comprehensive framework should cover and describes how Agent Spec addresses each of these requirements.
-Table 1: Challenges and objectives in agentic frameworks and solutions that we define with Agent Spec.
+Table 1 outlines key objectives that a comprehensive framework should cover and describes how Agent Spec addresses each of these requirements.
+Table 1: Addressing challenges and objectives in agentic frameworks through the Open Agent Specification.
 Topic Objective Open Agent Specification
-Agents
-Author and edit conversational solutions
-such as ReAct [15] style agents.
-Ability to define components that represent "resources"
-(with no impact on conversation flow routing) such as
-LLMs, memory, prompt templates, tools, other Agents
-which can be attached to other components’ properties.
-Flows
-Author and edit graph-style workflows
-such as multi-step business processes.
-Ability to define components that have a well-defined
-execution path such as a directed graph which may
-contain conditional routes or loops.
-Control flow routing
-Ability to define the flow of execution
-between components in the system,
-including conditional or looping flows.
-Explicit "relationships" (edges) for control flow
-supporting one-to-many and many-to-one transitions. The
-specific transition to be taken is determined at
-runtime for a given execution.
-I/O routing
-Ability to define which outputs are
-consumed by which inputs for fixed
-flow components. This includes
-conditional or looping flows, and outputs
-that are consumed in multiple places.
-Explicit "relationships" (edges) for data flow. Support
-for one-to-many and many-to-one transitions where the specific
-transition to be taken is determined at runtime for a
-given execution.
-Nesting/Composition
-Reuse components or solutions at any
-level of granularity in other solutions.
-Includes multi-agent composition.
-Definition of components that can be encapsulated and
-reused elsewhere. Established required properties for
-such components. Varying levels of visibility of
-exchanged messages between components.
-Orchestration of execution order and dependency either
-statically or dynamically.
-Referencing
-Ability to define a component instance
-once and refer to it in multiple places.
-"Reference" syntax to specify property values by
-referencing and reuse. Provision of components for
-defining an element and for using them.
-For example, an Agent component to define agent and
-an AgentNode in a Flow to use it in a certain place,
-or a Tool component to define a tool and a ToolNode
-in a Flow to use the tool in a specific place.
+Agents Author and edit conversational solutions such as ReAct [yao2023react] style agents. Define “resource” components (LLMs, memory, tools, or agents) attachable to other components’ properties.
+Flows Author and edit graph-style workflows such as multi-step business processes.
+Define components with well-defined execution paths (such as directed graphs), supporting conditional or looping routes.
+Control flow routing Ability to define the flow of execution between components in the system, including conditional or looping flows.
+Define control flow edges enabling one-to-many or many-to-one transitions where the specific transition to be taken is determined at runtime for a given execution.
+I/O routing Ability to define how outputs feed inputs for fixed, conditional, or looping flows, including reuse across steps.
+Define data flow edges supporting one-to-many or many-to-one transitions where the specific transition to be taken is determined at runtime for a given execution.
+Nesting & Composition Reuse components or solutions at any level of granularity in other solutions; includes multi-agent composition.
+Define components that can be encapsulated and reused; control message visibility, and orchestration of execution order and dependency either statically or dynamically.
+Referencing Ability to define a component instance once and refer to it in multiple places.
+Use reference syntax to specify property values and reuse defined components across multiple flows or tools.
+For example, define an Agent once and reference it via an AgentNode in a Flow.
 Moreover, different frameworks are usually parameterized and configured by different means.
 This makes porting solutions between frameworks a tedious and error-prone process, hindering collaboration and knowledge sharing among teams developing agentic applications.
-Consequently, a declarative, portable, and framework-agnostic representation of agentic designs is needed.
-For instance, ONNX [10] has revolutionized deep learning by providing a consistent way to port ML models across different frameworks (e.g., PyTorch [11], TensorFlow [13]).
-Similarly, Agent Spec aims to establish a unified representation for AI agents, enabling seamless interoperability and execution across diverse agentic frameworks.
-Just as ONNX allows models to be trained in one framework and executed in another, Agent Spec allows AI agents to be designed once and deployed across multiple platforms without reimplementation.
+To address these challenges, a declarative, shareable, portable, and framework-independent representation of agentic design is essential.
+Analogous to how ONNX [onnx] standardized ML model interchange across frameworks such as PyTorch [pytorch] and TensorFlow [tensorflow], Agent Spec aims to define a common interchange format for AI agents and workflows.
+Just as ONNX enables models trained in one framework to be executed in another without reimplementation, Agent Spec allows agents to be authored once and deployed consistently across multiple frameworks.
+2.3 Agent Spec within the Ecosystem
+Agent Spec streamlines the design of agentic assistants and workflows by providing a unified, framework-agnostic configuration language that defines agent patterns, components, and behaviors.
+It complements, rather than replacing, ongoing standardization efforts such as:
+  • Anthropic’s Model Context Protocol (MCP) [mcp]: Standardizes resource and data provision via client/server based RESTful APIs.
+  • Google’s Agent2Agent Protocol [a2a] and BeeAI’s Agent Communication Protocol (ACP) [acp]: Standardized APIs for distributed agent communication.
+  • The Agntcy initiative [agntcy] fosters open interoperability and governance standards, defining shared principles for agent discovery and cross-framework collaboration.
+While protocols like MCP and A2A standardize tool or resource provisioning as well as inter-agent communication, Agent Spec complements these efforts by enabling standardized configuration of components related to agentic system design and execution in general, as shown in Fig.
+1.
+Figure 1: Position of Agent Spec within the emerging agent-standardization stack.
+MCP governs resource and data provisioning, A2A/ACP specify inter-agent communication, and Agent Spec defines the declarative layer for agent behavior and execution semantics.
+3 Agent Spec: Objectives and Contributions
 3.1 Objectives of Open Agent Specification
-Agent Spec aims to define a common representation and act as a superset of capabilities for AI agents and workflows.
-Agent Spec graphs can be executed in different frameworks by means of runtimes that map Agent Spec components to the underlying framework’s primitives.
-Agent Spec, along with importers and exporters for various underlying frameworks, enables transparent portability of agent workflows.
-Agent Spec also provides methods to validate agent workflows before execution, ensuring they are syntactically and logically correct and executable.
-1.3 Key Benefits of Agent Spec
+Agent Spec defines a common representation that generalizes the capabilities for AI agents and workflows.
+Agent Spec graphs can be executed in different frameworks through runtimes that map its components to framework-specific primitives.
+Combined with importers and exporters for various frameworks, this enables transparent portability of agentic workflows.
+Agent Spec also provides validation mechanisms to ensure that workflows are syntactically and logically correct before execution, promoting consistency and reliability.
+3.2 Key Contributions of Agent Spec
+Agent Spec offers several key contributions that address current challenges in the development and deployment of agentic workflows.
+The following subsections highlight how Agent Spec improves portability, reusability, interoperability, and unification across diverse frameworks, ultimately enhancing the effectiveness and robustness of AI agent systems.
 3.2.1 Portability & Reusability
-  • Framework-Agnostic: Agent Spec abstracts agent definitions from specific implementations by providing a unified declarative representation, allowing agents to operate uniformly across multiple platforms (e.g., AutoGen, LangGraph, OCI Agents) without any need for reimplementation.
-  • Modular Design: Agent Spec’s component-based structure fosters reusability and extensibility, enabling enterprises to build complex agentic systems by assembling and configuring standardized components.
-  • Support for Complex Agentic Flows: Agent Spec facilitates the creation of intricate agentic workflows by supporting various modular components and reusable Agent Spec-defined patterns, enabling streamlined development of advanced AI agent systems.
+Framework-Agnostic: Agent Spec abstracts agent definitions from specific implementations by providing a unified declarative representation, allowing agents to operate uniformly across multiple platforms (e.g., AutoGen, LangGraph, WayFlow) without any need for reimplementation.
+Modular Design: Its component-based architecture promotes reusability and extensibility, allowing complex agentic systems to be composed from standardized primitives.
+Support for Complex Workflows: Agent Spec supports advanced workflows and modular design patterns, facilitating the seamless development of sophisticated AI agents.
 Portable Execution Graphs: Agent Spec enables execution graphs to be described in a way that is independent of the underlying runtime implementation, ensuring seamless portability across frameworks.
-Converters between the representations of other frameworks to Agent Spec would simplify the porting of previously developed agents to Agent Spec-compatible frameworks.
+Converters between the representations of other frameworks to Agent Spec simplifies the porting of previously developed agents to Agent Spec-compatible frameworks.
+A recent example is deprecation of AutoGen framework and corresponding migration to MS Agent Framework [autogenmigration].
 3.2.2 Interoperability & Compatibility
@@
 This simplifies development, deployment, and debugging across different environments.
-  • Robustness & Consistency: Agents defined in Agent Spec can be built on top of components that have been defined and validated within the Agent Spec specification, leading to higher reliability and predictability.
-  • Define Once, Run on Multiple Frameworks: Agentic systems leveraging Agent Spec can run on various frameworks through runtime adapters.
-Agent Spec’s open specification allows developers to implement these adapters to support any framework, enabling agents to run in optimized environments.
+Robustness & Consistency: Agents defined in Agent Spec can be built on top of components that have been defined and validated within the Agent Spec representation, leading to higher reliability and predictability.
+Define Once, Run on Multiple Frameworks: Agentic systems leveraging Agent Spec can run on various frameworks through runtime adapters, fostering open interoperability.
+This open specification allows developers to implement these adapters to support any framework, enabling agents to run in multiple environments.
 3.2.3 Unification across Frameworks
-  • Common Format: Agent Spec’s common format simplifies integration of agents and enables easier agent sharing across platforms.
-  • Knowledge Exchange: The unified declarative format that Agent Spec provides helps developers as well as maintainers of agents to efficiently exchange solutions and ideas.
-  • Conformance Test Suite: The Agent Spec conformance test suite offers a comprehensive set of tests that cover the various functionalities and components available within the specification, such as Agent, Flow, and ToolNode.
-It includes illustrative validation examples of Agent Spec YAML configurations and verifies that these configurations produce consistent execution outcomes and behaviors across multiple runtime environments.
+Common Format: Agent Spec’s common interchange format simplifies integration of agents and enables easier agent sharing across platforms.
+Knowledge Exchange: The unified declarative format that Agent Spec provides helps developers as well as maintainers of agents to efficiently exchange agent ideas and implementations.
+Conformance Test Suite: The Agent Spec conformance test suite offers a comprehensive set of tests that cover the various functionalities and components available within the specification, such as Agent, Flow, BranchingNode, MapNode, and ToolNode.
+It includes illustrative validation examples of Agent Spec configurations and verifies that these configurations produce consistent execution outcomes and behaviors across multiple runtime environments.
 This approach ensures interoperability and correctness of implementations supporting the Agent Spec specification.
-2 Open Agent Specification Design
-This section provides an overview of Agent Spec’s declarative model.
-It details how Agent Spec functions across different frameworks and outlines implementation specifics, including the SDK structure, supported programming languages, agent definition using the SDK, and the serialization/deserialization of agents and workflows.
-Agent Spec is intended to be a portable, platform-agnostic configuration language that allows agents and agentic systems to be described with sufficient fidelity.
-It defines the conceptual objects, called components, that compose agents in typical agentic systems, including the properties that determine the components’ configuration, and their respective semantics.
-Runtimes implement the Agent Spec components for execution with agentic frameworks or libraries.
-Agent Spec would be supported by SDKs in various languages (e.g., Python), enabling serialization and deserialization of Agents to YAML, or creating them from object representations with assurance of conformance to the specification.
-An outline of the components involved in the Agent Spec ecosystem is depicted in Fig. 1.
-Table 2: Agent Spec aims to do for AI agents what ONNX did for ML models
+4 Design Principles and Key Components
+This section provides an overview of the declarative model defined by Agent Spec.
+It explains how Agent Spec operates across multiple frameworks and outlines design details including SDK structure, supported programming languages, agent definitions using the SDK, and serialization and deserialization processes for agents and workflows.
+Figure 2 provides an overview of the Agent Spec ecosystem.
+Agent Spec is a portable, shareable, platform-agnostic configuration language that describes agents and agentic systems with sufficient fidelity for execution across runtimes.
+It defines the conceptual objects, called components, that make up agents in typical agentic systems, together with the properties that configure their behavior and semantics.
+Runtimes implement these Agent Spec components for execution within their own frameworks.
+Agent Spec is supported by SDKs in multiple languages (starting with Python), enabling serialization and deserialization of Agents to and from JSON (or YAML), or direct creation from object representations while maintaining full specification compliance.
+As shown in Table 2, Agent Spec provides for AI agents what ONNX provides for ML models.
+Figure 2: Agent Spec’s Design.
+4.1 Overview of Agent Specification
+Table 2: Agent Spec serves a similar unifying role for AI agents as ONNX does for ML models.
 Feature ONNX (ML Models) Agent Spec (AI Agents)
-Scope
-Unified representation of
-ML models, allowing portability
-across different deep learning
-frameworks.
-Unified representation of AI agents
-and workflows, enabling them to
-run across diverse agentic frameworks.
-Portability
-ONNX allows ML models to be
-trained in one framework (e.g.,
-PyTorch, TensorFlow) and
-executed in another.
-Agent Spec enables AI agents to be built in
-one framework (e.g., LangGraph [12], AutoGen [9],
-OCI Agents [7]) and run in another without modification.
-Unification
-Provides a common declarative
-format which is versioned as
-an Opset.
-Provides a common agent
-configuration format.
-Extensibility
-Supports various ML
-operations and optimizations
-Defines modular components for scalable agent systems.
-Figure 1: Agent Spec’s Design
-2.1 Overview of Agent Specification
+Scope Unified representation of ML models, allowing portability across different deep learning frameworks.
+Unified representation of AI agents and workflows, enabling them to run across diverse agent frameworks.
+Portability ONNX allows ML models to be trained in one framework (e.g., PyTorch, TensorFlow) and executed in another.
+Agent Spec enables AI agents to be built in one framework (e.g., LangGraph [LangGraph], AutoGen [autogen], crewAI [crewAI]) and run in another one.
+Unification Provides a common declarative format which is versioned as an Opset. Provides a common agent configuration format with versioning support.
+Extensibility Supports various ML operations and optimizations. Defines modular components for scalable agent systems.
 The Open Agent Specification defines structure and behavior of the conceptual building blocks, called components, that make up agents in typical agent-based systems.
-This structure can be serialized into a serialization language, like JSON or YAML.
-An overview of the language specification of Agent Spec follows, while the full specification can be accessed at https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html.
+The following sections describe each component in greater detail; the full specification can be found in [language_spec_25_4_1].
 4.2 Components
-In Agent Spec, components are the building blocks which cover the commonly used elements in agentic systems, such as (but not limited to) LLMs and tools.
+In Agent Spec, components are the building blocks which cover the commonly used elements in agentic systems, such as (but not limited to) Agents, LLMs, tools, etc.
 Describing these components with specific types enables:
-  • Type safety + Usage hints for GUIs
-    • Example: if a component’s property is expecting an LLM, only allow LLMs to be connected to it.
-    • Example: highlight valid connections from one component to others in a graphical programming interface for agents.
-  • Static analysis + validation
-    • Examples: if a component is a Flow, it must have a start node. If a component is an Agent, it must have an LLM.
-  • Ease of programmatic use
-    • Component families each have corresponding class definitions in the Agent Spec SDK.
-This allows consumers (agent execution environments, editing GUIs) to utilize concrete classes, rather than having to implicitly understand which type has which properties.
+  • Type safety and GUI guidance: For example, if a component property expects an LlmConfig, only compatible extensions can be connected, or valid connections can be highlighted to improve the agent-building experience in graphical interfaces.
+  • Static analysis and validation: For instance, if a component is a Flow, it must include a start node; if it is an Agent, it must contain an LLM.
+  • Ease of programmatic use: Component families have corresponding class definitions in the Agent Spec SDKs (see Section 5).
+This allows consumers (e.g., execution environments or GUI editors) to work with concrete classes directly, rather than inferring which properties belong to which types.
 4.2.1 Base Component
 The basic block in Agent Spec is a Component, which by itself can be used to describe any instance of any type, guaranteeing flexibility to adapt to future changes.
-Note that Agent Spec does not need to encapsulate the implementation code it describes; it simply needs to be able to express enough information to instantiate a uniquely identifiable object of a specific type with a specific set of property values.
+It is important to note that Agent Spec does not need to encapsulate the implementation code it describes; it simply needs to be able to express enough information to instantiate a uniquely identifiable object of a specific type with a specific set of property values.
 A metadata field contains all additional information that could be useful in different usages of the Agent Specification.
-For example, GUIs will require including some information about the components (e.g., position coordinates, size, color, notes, …), so that they will be able to visualize them properly even after an export-import process.
+For example, GUIs will require including some information about the components (e.g., position coordinates, size, color, notes), so that they will be able to visualize them properly even after an export-import process.
 Metadata is optional and set to null by default.
@@
 When a component references another component entity (for example as a property value), there is a simple symbolic syntax to accomplish this, "\$component_ref:" followed by the id of the other component.
-This type of relationship is applicable to both Agent components as well as Flow components (a fixed flow component may also use an LLM or memory component).
+"\$component_ref:{COMPONENT_ID}"
+This type of relationship is applicable to both Agent components as well as Flow components (a workflow component may also use an LLM or memory component).
 We do not need a separate object for this type of reference, as the reference is explicitly assigned to some property or parameter.
-1"\$component_ref:{COMPONENT_ID}"
 Input/output schemas
-Components might need some input data to perform their task and expose some output as result.
+Components might require input data to perform their task and produce output as result.
 These inputs and outputs must be declared and described in Agent Spec, so that users, developers and other components of the agent are aware of what is exposed as inputs and outputs respectively.
-Note that we do not add input/output schemas directly to the base Component class, as there are a few cases where they do not really apply (for example LLMConfig and edges, please see the next sections).
+Note that input/output schemas are not added directly to the base Component class, as there are a few cases where they do not apply (for example LLMConfig and edges; see the following sections for details).
 Input and output properties
-Input and output schemas are used to define which values the different components accept as input or provide as output. We call these values "properties".
-The term "properties" originates from JSON schema [8].
-In the Agent Spec specification, we rely on this widely accepted and consolidated format.
-For more information about JSON schema definition, please check the official website at https://json-schema.org/understanding-json-schema.
-In Agent Spec, a valid schema requires specific attributes:
-  • title: the name of the property
-  • type: the type of the property, following the type system described below
-Additionally, users can specify:
-  • default: the default value of this property
-  • description: a textual description for the property
-These are the minimal set of attributes (or annotations, as per JSON schema terminology) that we require to be supported for Agent Spec compatibility.
-Type system
-We rely on the typing system defined by the JSON schema format (see https://json-schema.org/understanding-json-schema/reference/type).
-At its core, JSON Schema defines the basic types available in most programming languages, though they may go by different names.
+Input and output schemas define the "values" that different components can accept as input or produce as output.
+These values are referred to as "properties" based on JSON schema standard [jsonschema].
 Inputs and outputs of nested components
-In case of nested components (for example using an Agent component inside a Flow, or a Flow inside another Flow, or just a step in a Flow), the wrapping component is supposed to expose a (sub)set of the inputs/outputs provided by the inner components together with additional inputs/outputs it might generate as shown in Fig.
-2.
-We require to replicate the JSON schema of each value in the inputs/outputs lists of every component that exposes it, hence, if a wrapper component exposes some inputs and outputs of some of its internal components, it will have to include them in its own list of inputs/outputs as well.
-This makes the specification of the components a bit more verbose, but clearer and more readable.
+In the case of nested components (for example, using an Agent component inside a Flow, or a Flow inside another Flow, or just a step in a Flow), the wrapping component is supposed to expose a (sub)set of the inputs/outputs provided by the inner components together with additional inputs/outputs it might generate as shown in Fig.
+3.
+Figure 3: Illustration of input/output exposure in nested components.
+Each value listed in the inputs and outputs must replicate its corresponding JSON schema in every component that exposes it.
+Consequently, if a wrapper component makes available certain inputs or outputs from its internal components, these must also be explicitly included in the wrapper’s own input/output lists.
+This makes the specification of the components more verbose, but transparent and more readable.
 Parsers of the specification are required to ensure the consistency of the input/output schemas defined in it.
-Figure 2: Inputs and outputs of nested components, as well as those of their wrapping component
-In the example above the wrapping Flow component (see sections below for more details about Flows) requires two inputs that coincide with two of the inputs exposed by the nested components and exposes one output among those exposed by the inner components in addition to one it computes itself.
+In the example above, the wrapping Flow component (see sections below for more details on Flows) requires two inputs that match those exposed by its nested components and produces one output derived from the inner components, along with an additional output computed internally.
 Validation of specified inputs/outputs
 Some input and output schemas are automatically generated by components based on their configuration.
-However, for the sake of clarity and readability of Agent Spec, we require that the descriptions of inputs and outputs are always reported in the Agent Specification.
-Consequently, when a serialized specification is read and imported by a runtime or an SDK, the inputs and outputs configuration of each component must be validated against the one generated by its configuration (if any).
+However, for the sake of clarity and readability of Agent Spec, all input and output descriptions must be explicitly included in the Agent Specification.
+Consequently, when a serialized specification is loaded by a runtime or SDK, the inputs and outputs configuration of each component must be validated against those automatically generated from its configuration (if applicable).
 Specifying inputs through placeholders
-Some components might infer inputs from special parts of their attributes.
-For example, some nodes (e.g., the LLMNode) extract inputs from their configuration attributes, so that their values can be adapted based on previous computations We let users define placeholders in some attributes, by simply wrapping the name of the property among a double set of curly brackets.
-The node automatically infers the inputs by interpreting those configurations and the actual value of the input/output will be replaced at execution time.
+Some components can infer inputs from specific parts of their configurations.
+For example, certain nodes (e.g., the LLMNode) extract inputs from their configuration attributes, allowing their values to be dynamically adapted based on previous computations.
+Users can define placeholders within these attributes by enclosing the name of a property in double curly brackets.
+The node automatically infers the inputs by interpreting those configurations and the actual value of the input/output will be substituted at execution time.
 Whether an attribute accepts placeholders depends on the definition of the component itself.
@@
 The Agent is the top-level component that holds shared resources such as conversation memory and tools.
-It also represents the entry point for interactions with the agentic system.
-Its purpose is to fill out the values for all the properties defined in the outputs attribute, with or without any interaction with the user.
-It is important to have a separate definition for Agents as components, so that the same Agent component can be reused several times in a Flow or as part of a more complex agent design, without replicating its definition multiple times.
+It also serves as the entry point for interactions with the agentic system.
+Its primary goal is to populate the values of all properties defined in its output attribute, either through user interaction or autonomous computation.
+Defining Agents as standalone components is essential, as it enables the same Agent component to be reused across multiple Flows or within more complex agent architectures, without the need to duplicate its definition.
 4.2.3 LLM
-Large Language Models (LLM) are used in agents as generative models.
-To embed them into the agent, they need to be configurable flexibly with configuration parameters like the connection details, model id, the generation parameters, and more.
-By default, only the generation parameters used when prompting the LLM must be specified. Specific extensions of LLMConfig for the most common model providers can be derived.
-Indeed, this is not limited to locally hosted models; in principle, it also enables the configuration of API endpoints such as those provided by OpenAI, Anthropic, Mistral, and others.
+Large Language Models (LLMs) are used within agents as generative components.
+To integrate them effectively, they must be flexibly configurable through parameters such as connection details, model identifiers, generation settings, and other options.
+By default, only the generation parameters required for prompting the LLM need to be specified.
+Specific extensions of LLMConfig can be derived to support the most common model providers.
+This approach is not limited to locally hosted models; it also supports the configuration of external API endpoints, such as those offered by OpenAI, Anthropic, Mistral, and others.
 4.2.4 Tool
 A tool is a procedural function or a Flow that can be made available to an Agent component to execute and perform tasks.
-The Agent component can decide to call a tool based on its signature and description.
-In a fixed Flow context, a node would need to be configured to call a specific tool.
-In Agent Spec, we differentiate tools according to where the actual functionality is executed or run and specify three corresponding types:
-  • ServerTools are executed in the same runtime environment where the agent is being executed.
-The definitions of these tools therefore must be available to the agent’s environment.
-It is expected that this type of tool will be limited in number, and generally in functionality.
-  • ClientTools are not executed by the runtime. The client environment must execute the tool and provide the results back to the runtime (like OpenAI’s function calling model).
-  • RemoteTools are run in an external environment and are triggered by an RPC or REST call from the agent runtime.
-Agent Spec is not supposed to provide an implementation of how these types of tools should work but provides their representation, so that they can be correctly interpreted and ported across platforms and languages.
-In principle a Tool is a function that is called providing some parameters (i.e., inputs), performs some transformation, and returns the result (i.e., outputs).
-The function has a name, and a description that can be used by an LLM for context when it is relevant to perform a task.
-Therefore, we define Tool as an extension of ComponentWithIO.
-Tools may specify multiple outputs.
-In this case, the expected return value of the tool is a dictionary where each entry corresponds to an element specified in the output.
-The key of the dictionary entry must match the name of the property specified in the output.
+The Agent can decide which tool to invoke based on the tool’s signature and description.
+In a Flow context, a node must be explicitly configured to call a specific tool.
+A novelty in Agent Spec is the classification of tools according to where their actual functionality is executed. Four corresponding types are defined:
+  • ServerTools are executed within the same runtime environment as the agent itself.
+The definitions of these tools therefore must be available in the agent’s environment.
+Such tools are typically limited in number and functionality.
+  • ClientTools are not executed by the runtime.
+Instead, they run in the client environment, which is responsible for executing the tool and returning the results to the runtime (similar to OpenAI’s function calling model).
+  • RemoteTools are executed in an external environment and triggered through an RPC or REST call initiated by the agent runtime.
+  • MCPTools enable execution of tools provided by a MCP server in accordance with the MCP specification [mcp].
+Agent Spec does not prescribe how these tool types should be implemented.
+Rather, it provides a representation format that allows them to be correctly interpreted and exchanged across different platforms and programming languages.
+Conceptually, a Tool is a function that accepts parameters (inputs), performs a transformation, and returns a result (outputs).
+Each tool has a name and a description, which can provide context to an LLM when deciding which tool to use for a given task.
+Consequently, a Tool is defined as an extension of ComponentWithIO.
+Tools may produce multiple outputs.
+In such cases, the expected return value is a dictionary, where each key corresponds to the name of an output property defined in the tool’s schema.
 The runtime should parse the dictionary to extract the different outputs and bind them correctly.
-While ServerTool and ClientTool do not require specific additional parameters, the
-RemoteTool requires including also the details needed to invoke the remote tool call.
-An important security consideration for Agent Spec is the exclusion of arbitrary code from the agent specification.
-The specification is limited to a comprehensive description of the tool—including its attributes, inputs, outputs, and other relevant details—but does not contain any executable code.
+While ServerTool and ClientTool do not require additional parameters beyond their standard definition, a
+RemoteTool must include the necessary details for invoking the remote call.
+Finally, an important security consideration in Agent Spec is the exclusion of arbitrary code from the specification.
+The Agent Specification only provides a structured description of each tool, including its attributes, inputs, outputs, and metadata but does not contain any executable code.
 4.2.5 Flow
 Flows (or graphs) are directed, potentially cyclic workflows.
-They can be thought of as "subroutines" that encapsulate consistently repeatable processes.
-As such, they provide more determinism and reliability compared to puristic Agent components at the cost of flexibility.
-Flows are executable as tools by Agent components.
-Each Flow requires 0 or more inputs (any input edges expressed on the starting node of the graph) and may produce 0 or more outputs (any output edges expressed by the terminal node of the graph - note that a graph can have more than one terminal node, in the case of branching logic).
-As such, the I/O surface of a Flow is directly mutually exchangeable with that of a tool.
-Flow objects specify the the directed graph’s entry point, nodes, and edges. Components inherent in the Flow are described in subsequent sections.
-A Flow exposes the set of outputs that are available in all EndNode objects, and all inputs that are defined in its StartNode.
+They can be regarded as "subroutines" that encapsulate structured sequences of operations.
+As such, Flows offer greater determinism and reliability compared to purely agent-driven components at the cost of reduced flexibility.
+Each Flow may require zero or more inputs (corresponding to the inputs expressed on the starting node of the graph) and may produce zero or more outputs (represented by the terminal node of the graph - note that a graph might contain multiple terminal nodes in cases involving branching logic).
+Flow objects specify the directed graph’s entry point, its nodes, and the edges connecting them. The individual components that compose a Flow are described in detail in the following sections.
+A Flow exposes all inputs defined in its StartNode and all outputs that are available across its EndNode objects.
+Conversation
+At the core of a conversational agent’s execution lies the conversation itself.
+The conversation is implicitly propagated across all components within the Flow and maintains the collection of messages generated throughout execution.
+Each node in a Flow, as well as each Agent, might append messages to this conversation as needed.
+Every message contains, at minimum, the following attributes: role (e.g., system, agent, user), content, sender, and recipient.
+Input/Output Data Space
+Alongside the conversation, flows have another data space composed of inputs and outputs generated by nodes.
+When a node generates an output, this is exposed by the node itself, so that other nodes can consume it as input, if needed.
+The input-output connections, called data flow, can be defined through explicit data relationships, or through implicit, name-based property matching, as depicted in the following sections.
 Relationships / Edges
-For Flows, Agent Spec defines separate relationships (edges) for both properties and control flow.
-This allows Agent Spec to unambiguously handle patterns such as conditional branches and loops, which are challenging to support with data-only flow control.
-Relations are only applicable for Flows. All relations express a transition from some source node to some target node.
+Within Flows, Agent Spec distinguishes between two categories of relationships (or edges): those governing data flow and those governing control flow.
+This distinction enables Agent Spec to unambiguously represent complex execution patterns such as conditional branches and loops, which are difficult to model using data flow alone.
+Relationships are defined exclusively within the context of Flows. Each relationship expresses a directed transition from a source node to a target node.
 4.2.6 ControlFlowEdge
-A control relationship defines a potential transition from one branch of a node to another.
-The actual transition to be taken during a given execution is defined by the implementation of a specific node.
-The from_branch attribute of a ControlFlowEdge is set to null by default, which means that it will connect the default branch (i.e., “next” ) of the source node.
+A control flow relationship specifies a potential transition from one branch of a node to another.
+The actual transition followed during execution is determined by the implementation of the corresponding node.
+The attribute from_branch of a ControlFlowEdge is set to null by default, indicating that the edge connects to the node’s default branch (i.e., next).
 4.2.7 DataFlowEdge
-Considering an I/O system, a component id alone is not sufficient to identify a data relationship.
-To this purpose, a data relationship is needed to define what output property value should be used to fill an input property.
-A component may take multiple input parameters and/or produce multiple output values.
-We must be able to determine what value or reference is mapped to each input and where each output is used.
-Destinations (Inputs of the destination node) can be connected to outputs of another node or to static values.
-Sources (Outputs of the source node) can be connected to inputs of another node or left unconnected.
+In an input/output system, a component identifier alone is insufficient to fully specify a data relationship.
+Therefore, Agent Spec defines explicit data flow relationships to describe how the output property of one node is mapped to the input property of another.
+A component might accept multiple input parameters and/or produce multiple output values.
+Hence, it is necessary to define which output or reference is associated with each input, and conversely, how each output is consumed.
+Destinations (the inputs of the target node) can be connected to outputs from another node or assigned to static values.
+Conversely, sources (the outputs of the source node) may be connected to inputs of subsequent nodes or left unconnected if unused.
 Connecting multiple data edges
-(a) Control flow with two branches that converge into the same node with a property filled based on the branch taken
+(a) Control flow with two branches that converge into the same node with a property filled based on the branch taken.
 (b) Control flow with a self-loop and a property updated by consequent execution of the same node
-Figure 3: Possible connections formed by ControlFlowEdges and DataFlowEdges. ControlFlowEdges are shown as solid lines, DataFlowEdges with dotted lines.
-In the definition of a Flow, it is allowed to connect two or more different data outputs to the same input as shown in Fig.
-3.
-The most recently executed node with an output connected to the input takes precedence.
-In other words, the behavior is like having the input exposed as a public variable and every node that has an output connected to that input updates its value.
-For the control flow, instead, edges just define which are the allowed "next step" transitions: multiple connections from the same outgoing control flow branch are not allowed, and they do not involve any ambiguity (note that we do not enable parallelism through edges).
-Optionality of data relationships
-Some frameworks do not require specifying the data flow explicitly as they assume that all the properties are publicly exposed in a shared "variable space" that any component can access, and the read/write strategy is simply name based.
+Figure 4: Control- and data-flow relationships in Agent Spec. Solid lines denote control-flow transitions, while dotted lines indicate data propagation.
+In the definition of a Flow, it is permissible to connect two or more distinct data outputs to the same input as shown in Fig.
+4.
+In such cases, the value produced by the most recently executed node among those connected to the input takes precedence.
+Conceptually, this behavior is analogous to exposing the input as a public variable, where each connected node can update its value upon execution.
+In contrast, control flow edges define the set of permitted next-step transitions.
+Multiple outgoing connections from the same control flow branch are not permissible, ensuring unambiguous execution order.
+It should also be noted that parallelism is not enabled through control flow edges.
+Optionality of data flow relationships
+Some frameworks do not require the explicit specification of data flow, as they assume that all properties are publicly exposed within a shared "variable space" accessible to any component.
+In such systems, read and write operations are determined purely by name-based resolution.
 In practice, this means that:
-  • When a component has an input with a specific name, it will access the public variable space to read the value of the variable with that name.
-  • When a component has an output with a specific name, it will write (or overwrite if it already exists) into the public variable space the variable with that name.
-We let users adopt this type of I/O system by setting the data_flow_connections parameter of the Flow to None.
-In this case, the name-based approach explained above will be adopted, by using the shared variable space described above as the public space where values are published.
-Note that this name-based approach can be expressed by explicitly defining data-flow connections (but the opposite is not possible) by connecting all the outputs to the inputs with the same name, following the control flow connections in the Flow to account for overwrites.
-Flow state
-We define as "flow state" the set of all values that describe the current stage of execution within the flow (the point of execution reached).
-Users have the ability to specify which set of values within the flow state are publicly exposed to all nodes.
-These values are treated differently compared to values that are exposed through inputs and outputs of nodes:
-  • They are always available in every node and do not need to be connected by a DataFlowEdge.
-  • If a component writes an output which name matches a flow state value, the respective value in the flow state is overwritten.
-Conversation
-At the core of the execution of a conversational agent, there is the conversation itself.
-The conversation is implicitly passed to all the components throughout the flow.
-It contains the messages being produced by the different nodes: each node in a flow and each agent can append messages to the conversation if needed.
-Messages have, at least, a type (e.g., system, agent, user), content, the sender, and the recipient.
+  • When a component defines an input with a given name, it retrieves the value of the variable with the same name from the shared variable space.
+  • When a component defines an output with a given name, it writes (or overwrites, if already present) the corresponding variable in the shared variable space.
+Users can adopt this type of I/O mechanism in Agent Spec by setting the data_flow_connections parameter of a Flow to None.
+In this configuration, the system follows the name-based approach described above, using the shared variable space as the public medium for publishing and accessing values.
+It is worth noting that this name-based strategy can be expressed equivalently by explicitly defining the corresponding data flow connections by connecting all outputs and inputs that share the same name and following the control flow to account for overwrites.
+However, the converse is not true: an explicitly defined data flow configuration cannot always be reduced to a purely name-based representation.
 4.2.8 Node
-A Node is a vertex in a Flow and is derived from ComponentWithIO.
-Agent Spec provides a library of nodes designed to streamline workflow orchestration, enabling rapid development and seamless integration for a wide range of applications:
-  • LLMNode: uses a LLM to generate some text from a prompt.
-  • APINode: performs an API call with the given configuration and inputs.
-  • AgentNode: runs a multi-round conversation with an Agent component, used to better structure agents and easily reuse Agent components.
-  • FlowNode: runs a Flow, used to better structure agents and easily reuse Flows.
-  • MapNode: performs a map-reduce operation on a given input collection, applying a specified Flow to each element in the collection.
-  • StartNode: entry point of a Flow.
-  • EndNode: exit point of a Flow.
-  • BranchingNode: allows conditional branching based on the value of an input.
-  • ToolNode: executes a tool.
-A more detailed description of each node type is given in https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html#standard-library-of-nodes-to-use-in-flows.
+A Node represents a vertex within a Flow and is derived from the ComponentWithIO class.
+Agent Spec provides a standard library of node types designed to facilitate workflow orchestration, enabling efficient development and seamless integration across a wide range of applications.
+The main node types include:
+  • LLMNode: Utilizes a large language model (LLM) to generate text from a given prompt.
+  • APINode: Performs an API call based on the specified configuration and inputs.
+  • AgentNode: Executes a multi-turn interaction with an Agent component, supporting modular agent design and reuse.
+  • FlowNode: Executes a Flow, enabling hierarchical structuring and reuse of Flows within more complex architectures.
+  • MapNode: Performs a map-reduce operation over an input collection, applying a specified Flow to each element in the collection.
+  • StartNode: Defines the entry point of a Flow.
+  • EndNode: Defines the termination point of a Flow.
+  • BranchingNode: Enables conditional branching based on the value of a specified input.
+  • ToolNode: Invokes and executes a defined Tool.
+  • InputMessageNode: Interrupts temporarily the execution to retrieve user input.
+  • OutputMessageNode: Appends an agent message to the conversation.
+A comprehensive description of each node type is provided in the Agent Spec documentation, available at [language_spec_25_4_1].
 4.3 Open Agent Specification Model
-The objective of Agent Spec is to be agnostic with respect to any agentic framework and the programming language used to build and execute an agent.
-For this reason, an agent represented with Agent Spec should be exportable to, and importable from a representation that is specified in a common serialization language.
-To this end, Agent Spec adopts JSON as the designated language for the serialization of its components.
-All building blocks of Agent Spec are designed with the goal of being trivially serializable.
-It follows that the Agent Specification of a component can be obtained by serializing its definition according to the JSON specification.
-In particular, the serialization of a component includes its attributes, plus an additional field called component_type that will be used at deserialization time to infer the correct component type to deserialize.
-Note that, as defined in the Agent Spec language, components can be declared only once in the representation and then referenced according to the rules described in the Symbolic references section of this document.
-A few examples of component JSON serializations are given at https://oracle.github.io/agent-spec/agentspec/language_spec_25_4_1.html#language-examples.
+The primary objective of Agent Spec is to remain agnostic with respect to any agentic framework and the programming language used to build and execute an agent.
+Consequently, an agent represented in Agent Spec should be exportable to, and importable from a common, framework-independent serialization format.
+To achieve this, Agent Spec adopts JSON as the designated language for the serialization of its components.
+All building blocks within Agent Spec are explicitly designed to be trivially serializable.
+Thus, the Agent Specification of any component can be directly obtained by serializing its definition according to the JSON standard.
+Each serialized component includes its set of attributes, along with an additional field, component_type, which is used during deserialization to infer the appropriate component type.
+As defined by the Open Agent Specification, each component may be declared only once within a representation and subsequently referenced symbolically, following the rules outlined in the Symbolic References section of this article.
+Illustrative examples of component serializations in JSON format are available at [language_spec_25_4_1].
 4.4 How Agent Spec Works Across Frameworks
-Agent Spec serves as an abstraction layer above framework-specific implementations, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints.
-Hence, the goal of Agent Spec is to define behavioral patterns for the different components that should be implemented according to the specificities of the different agentic frameworks.
-An Agent Spec runtime for a specific framework should be able to read the Agent Specification of a component (e.g., a Flow or an Agent), and then execute its logic, adhering to the behavior described in the Agent Spec.
-As an alternative, if the agentic framework - used to implement the runtime - offers a serialization mechanism, it is also possible to build a translation script between the specification of an agent and the equivalent framework specific representation, then execute the runtime on the latter.
+The primary objective of Agent Spec is to define behavioral patterns for the various components that can be implemented according to the specific characteristics of different agentic frameworks.
+A runtime environment developed for a particular framework should be capable of reading the serialized configuration of a component (such as a Flow or an Agent) and executing its logic in accordance with the behaviors prescribed by Agent Spec.
+Every Agent Spec configuration includes a top-level agentspec_version property specifying the specification version it follows.
+Alternatively, if the agentic framework used to implement the runtime provides its own serialization mechanism, a translation layer could be introduced.
+This layer can convert an Agent Spec configuration into the equivalent framework-specific representation, allowing the runtime to execute the agent directly using the translated format.
 4.5 Implementation Details
-2.5.1 Agent Spec SDKs
-To facilitate the process of building programmatically framework-agnostic agents, Agent Spec SDK packages can be conveniently implemented. These SDKs should provide two functionalities:
-  • Building Agent Spec component abstractions through the implementation of the relevant interfaces, in conformity with the Agent Spec.
-  • Import and export these abstractions from and to their respective serialized version in JSON format.
-As part of the Agent Spec effort, we provide Agent Spec’s Python SDK that we call PyAgentSpec.
-PyAgentSpec allows building Agent Spec-compliant agents in python.
-Users can create their own assistants using component classes that replicate the interface and behavior defined by Agent Spec, which can finally be exported in JSON format.
-Here is an example of how the definition of a simple agentic Flow prompting an LLM with a custom user input and returning the LLM’s output would look like in PyAgentSpec.
-1from pyagentspec.property import Property
-2from pyagentspec.flows.flow import Flow
-3from pyagentspec.flows.edges import ControlFlowEdge, DataFlowEdge
-4from pyagentspec.flows.nodes import LlmNode, StartNode, EndNode
-5from pyagentspec.llms import VllmConfig
-6llm_config = VllmConfig(
-7 name="<Your Model Name Here>",
-8 url="<url.of.your.llm.deployment:port>",
-9 model_id="<provider/model-identifier>",
-10)
-11prompt_property = Property(
-12 json_schema={"title": "prompt", "type": "string"}
-13)
-14llm_output_property = Property(
-15 json_schema={"title": "llm_output", "type": "string"}
-16)
-17start_node = StartNode(name="start", inputs=[prompt_property])
-18end_node = EndNode(name="end", outputs=[llm_output_property])
-19llm_node = LlmNode(
-20 name="simple llm node",
-21 llm_config=llm_config,
-22 prompt_template="{{prompt}}",
-23 inputs=[prompt_property],
-24 outputs=[llm_output_property],
-25)
-26flow = Flow(
-27 name="Simple prompting flow",
-28 start_node=start_node,
-29 nodes=[start_node, llm_node, end_node],
-30 control_flow_connections=[
-31 ControlFlowEdge(name="start_to_llm", from_node=start_node, to_node=llm_node),
-32 ControlFlowEdge(name="llm_to_end", from_node=llm_node, to_node=end_node),
-33 ],
-34 data_flow_connections=[
-35 DataFlowEdge(
-36 name="prompt_edge",
-37 source_node=start_node,
-38 source_output="prompt",
-39 destination_node=llm_node,
-40 destination_input="prompt",
-41 ),
-42 DataFlowEdge(
-43 name="llm_output_edge",
-44 source_node=llm_node,
-45 source_output="llm_output",
-46 destination_node=end_node,
-47 destination_input="llm_output"
-48 ),
-49 ],
-50)
-To export the Flow that was just built in JSON format, the code is the following.
-1serializer = AgentSpecSerializer()
-2serialized_flow = serializer.to_json(flow)
-The full documentation of PyAgentSpec, including installation instructions, APIs documentation and examples, is available at https://oracle.github.io/agent-spec/index.html.
-2.5.2 Agent Spec SDK Plugins
-The PyAgentSpec SDK also supports the concept of plugins, because we expect that teams using Agent Spec may want to create new component types to be included in their configurations.
-To this end, a plugin can enable newer components in addition to the standard components of Agent Spec, simply by specifying the list of component types that the plugin includes and by providing the serialization and deserialization logic for the plugin.
-Such plugins can be implemented for components not yet included in the first version of Agent Spec, such as multi-agent patterns, or planning and memory components.
-See below the abstract interface to be implemented for a serialization plugin (the deserialization interface follows the same logic).
-1from abc import ABC, abstractmethod
-2from typing import Any, Dict, List
-3
-4from pyagentspec.component import Component
-5from pyagentspec.serialization.serializationcontext import SerializationContext
-6
-7class ComponentSerializationPlugin(ABC):
-8 """Base class for Component serialization plugins."""
-9
-10 @property
-11 @abstractmethod
-12 def plugin_name(self) -> str:
-13 """Return the plugin name."""
-14 pass
-15
-16 @property
-17 @abstractmethod
-18 def plugin_version(self) -> str:
-19 """Return the plugin version."""
-20 pass
-21
-22 @abstractmethod
-23 def supported_component_types(self) -> List[str]:
-24 """Indicate what component types the plugin supports."""
-25 pass
-26
-27 @abstractmethod
-28 def serialize(
-29 self, component: Component, serialization_context: SerializationContext
-30 ) -> Dict[str, Any]:
-31 """Serialize a component that the plugin should support."""
-32 pass
-2.5.3 Agent Spec Runtime Adapters
-A runtime implementation makes Agent Spec components abstractions runnable by implementing the behavioral logic described in the Agent Spec language.
-As previously discussed, there are several ways to implement runtimes for Agent Spec.
-Ideally, a runtime implementation is an agentic framework that natively supports Agent Spec: it should read an Agent Spec JSON specification and transform it directly into an equivalent component that can be executed.
-As an example of an Agent Spec runtime implementation, we provide a package called WayFlow, which is a powerful, intuitive Python library for building sophisticated AI-powered assistants.
-It includes a library of plan steps to streamline the creation of AI-powered assistants, supports reusability, and it is ideal for rapid development.
-The WayFlow framework provides a serialization and deserialization interface to import and export Agent components according to the Agent Spec.
-It is a reference framework for Agent Spec and our runtime of reference, with native support for all Agent Spec Agents and Flows.
-When an agentic framework does not allow reading Agent Specifications directly, a translation layer that transforms the Agent Specifications into a framework-specific representation could be adopted.
-This layer, which we can call Agent Spec Runtime adapter, loads an Agent Specification and transforms it into an equivalent component of the specific agentic framework.
-Runtime adapters enable existing agentic frameworks that do not support Agent Spec natively executing Agent Specifications.
-For example, we also provide runtime adapters for some of the most popular agentic frameworks, including, but not limited to, LangGraph and Microsoft AutoGen.
-These adapters load Agent Specifications taking advantage of the Agent Spec Python SDK, and they transform the Agent Spec components into the respective versions defined in the agentic framework of interest.
-2.6 Performance Evaluation & Benchmarks
-Agent Spec enables key aspects of productizing agent-driven solutions, including comparison and evaluation of agent task performance across different executing frameworks and models, as well as optimizations (including costs) of execution for underlying hardware and frameworks.
-By covering commonly accepted concepts in agentic systems, such as workflows and agents, Agent Spec allows developers to map these concepts to various execution frameworks, facilitating comparison of framework specific task performance.
-Compute optimization is also possible, as Agent Spec enables hardware-optimized and framework-optimized implementation of agentic design patterns.
-2.6.1 Portability Analysis
-Agentic AI is rapidly evolving, and a unified benchmark for assessing agent reliability, particularly in enterprise settings, is currently lacking.
-Future updates for Agent Spec report will include empirical results demonstrating the reliability of Agent Spec-defined agents across different frameworks.
-https://oracle.github.io/agent-spec/howtoguides/howto_execute_agentspec_across_frameworks.html provides an example of how an agent defined with Agent Spec can be run in different frameworks.
-In this Retrieval Augmented Generation case study, an agent utilizing the same tools is instantiated in AutoGen, LangGraph, and WayFlow.
-In a practical setting, a specific mapping from Agent Spec to each framework would be implemented.
-2.6.2 Use Cases & Real-World Applications
-A common use case of Agent Spec is the porting of agent-based assistants from one framework to another.
-Let’s take as an example a team that implemented agent-based workflows using AutoGen.
-As business and process needs to change, evaluations are conducted to consider alternative frameworks.
-Due to AutoGen’s specific architecture, migrating to another agentic framework proves to be complex with significant rewriting overhead.
-In this example, Agent Spec enables the definition of business unit-specific agentic systems using a declarative representation.
-This allows for conversion to specific framework formats or flexible execution across different agentic frameworks with appropriate adapters, enabling straightforward comparison of task performance and runtime cost depending on the execution framework.
-Consequently, utilizing Agent Spec throughout the agent design and deployment process reduces complexity, costs, implementation effort, and potential inconsistencies.
-Another use case involves developing and deploying Agents across varied compute environments, such as cloud platforms and on-premise datacenters.
-In the cloud, Agents may run virtualized or serverless as part of a scalable system design.
-In an on-premise scenario, they may run within an application server.
-Agent Spec’s framework and runtime agnosticism enables development of the Agent in a local development environment, followed by flexible deployment across either of these scenarios.
-A different compelling real-world application involves advanced multi-agent systems in the domain of financial crime compliance.
-By integrating generative AI agents, it enables to automate key investigative processes - supplementing human investigators with narrative-driven analyses, uncovering critical insights, and recommending evidence-based actions.
-In this context, a comprehensive Anti-Money Laundering platform may comprise specialized agents tasked with sanctions screening, transaction monitoring for suspicious activity, and the generation of enhanced due diligence reports.
-These sub-agents are often developed by independent teams using a variety of frameworks and technologies, and may need to be deployed across heterogeneous environments, including cloud, on-premises, or even within database systems.
-Leveraging Agent Spec’s framework-agnostic and declarative approach, development teams can seamlessly integrate and orchestrate these diverse agent workflows, reducing integration overhead and enabling consistent, scalable deployment.
-This holistic and automated strategy minimizes manual effort, decreases operational risk, and empowers financial institutions to respond more rapidly and consistently to evolving financial crime patterns, regardless of the underlying deployment environment.
-2.7 Positioning in the Ecosystem
-Agent Spec aims to streamline the architecture and design of agentic assistants and workflows, serving as a unified representation that abstracts away implementation details of specific agentic frameworks.
-It is a portable configuration language that describes agentic design patterns, components and expected behavior.
-However, Agent Spec is not the only effort aimed at unifying the different parts that compose a common agentic ecosystem.
-  • The Model Context Protocol (MCP) [1], introduced by Anthropic, standardizes resource and data provision to agents. Resources are made available with a client/server based RESTful API.
-  • Google’s Agent2Agent Protocol [6] and BeeAi’s Agent Communication Protocol (ACP) [3] propose standardized APIs for distributed agent communication.
-  • Nvidia recently announced NeMo Agent toolkit which is a framework-independent library that treats “every agent, tool, and agentic workflow as a function call—enabling composability between these agents, tools, and workflows”[5].
-The idea is to compose existing agentic solutions and resources, potentially written using different frameworks, into new systems.
-While protocols like MCP and A2A standardize tool or resource provisioning as well as inter-agent communication, Agent Spec complements these efforts by enabling standardized configuration of components related to agentic system design and execution in general, as shown in Fig.
-4.
-Hence, Agent Spec provides strong synergies with other standardization efforts through its abstract definition of agentic system design.
-In the context of ongoing unification efforts within the agentic system landscape, Agent Spec aims to serve as the "common foundation" that connects these initiatives, enhancing their effectiveness and fostering a more cohesive ecosystem.
-Support for the aforementioned protocols may be added in future versions of Agent Spec as new components.
-Figure 4: Agent Spec complements other standardizations, such as MCP or A2A
-2.8 Future Roadmap & Enhancements
-2.8.1 Upcoming Features
-Expanding the Open Agent Specification language We are looking forward to extending Agent Spec with new concepts and components in the upcoming versions of Agent Spec. A few notable ones include:
-  • Memory
-  • Planning
-  • Datastores
-  • Support for remote agents (e.g., A2A agents)
-  • Support for new types of tools (e.g., MCP tools)
-Extending runtimes support
-We want to foster the implementation of Agent Spec runtime adapters for the popular agentic frameworks in the vibrant AI community.
-At the same time, we aim to offer a conformance test suite for runtimes to ensure consistent behavior in accordance with the Agent Specification.
-Improving user experience
-Even though Agent Spec SDKs simplify the creation and validation of Agent Specifications, creating agents by writing code might be complex and error-prone.
-We are working on a Drag&Drop UI that will allow users to build their own agents in a visual interface and then export the Agent Specification to be used in the framework of their choice via a runtime adapter.
-2.8.2 Community & Ecosystem Growth
-We strongly believe Agent Spec will contribute to streamlining agent development and open-sourcing its specification to encourage community contributions.
-This will foster a rich ecosystem of interchangeable and reusable designs and tools for agentic AI development.
-Integration of Agent Spec by other third-party open-source projects as an agent design language and specification will benefit the open-source community, which is currently evolving and lacks established common unified representation.
-In summary, the key aspects of impact on community growth are:
-  • Design patterns that can be easily shared & reused to collaboratively build increasingly complex solutions.
-  • Enable open exchange of building blocks and agentic patterns.
-  • Encourage contributions by the open-source community.
-To help guide the evolution of the Agent Specification in alignment with its core design principles, we will establish a steering committee.
-Similar to ONNX principles, this committee will work closely with the community to review proposed contributions, foster open discussion, and publish versioned updates.
+To facilitate the programmatic construction of framework-agnostic agents, Agent Spec supports the development of dedicated Software Development Kits (SDKs).
+Our reference SDK, PyAgentSpec, also includes a plugin architecture that allows teams adopting Agent Spec to define and integrate new component types within their configurations as needed.
+A runtime adapter implementation makes Agent Spec component abstractions executable by providing the behavioral logic defined in the Agent Spec language.
+Beyond simple translation, these runtime adapters act as “compiler” for Agent Spec, converting its declarative specifications into executables for target frameworks and bridges high-level declarative design with low-level execution semantics.
+Please see Appendix A for an example of an agentic flow in PyAgentSpec and Appendix B for details on the reference Agent Spec SDK, PyAgentSpec, the plugin architecture and runtime adapters.
 5 Evaluation & Benchmarking
@@
 5.1 Evaluation methodology
-We evaluate Agent Spec on three popular benchmarks: SimpleQA Verified [4], $\tau^{2}$-Bench [2] and BIRD-SQL [14].
+We evaluate Agent Spec on three popular benchmarks: SimpleQA Verified [haas2025simpleqaverified], $\tau^{2}$-Bench [barres2025tau2] and BIRD-SQL [li2024can].
 For each benchmark, we select one or more Agent Spec JSON configurations and execute the same specification across four framework runtimes: AutoGen, CrewAI, LangGraph, and WayFlow.
@@
 SimpleQA Verified.
-We evaluate on the full set of 1,000 questions using two representative assistants: (i) a ReAct-style agent equipped with a web-browsing tool, and (ii) a flow-based agentic RAG solution that applies question decomposition, web-browsing, and self-reflection to answer complex user queries [^1] Note that the browse tool makes the task [4] easier, but it is acceptable since we are demonstrating Agent Spec’s functional effectiveness instead of competing on the benchmark leaderboard..
+We evaluate on the full set of 1,000 questions using two representative assistants: (i) a ReAct-style agent equipped with a web-browsing tool, and (ii) a flow-based agentic RAG solution  [^3] Details about the flow is provided in Appendix D.1.
+that applies question decomposition, web-browsing, and self-reflection to answer complex user queries [^4] Note that the browse tool makes the task [haas2025simpleqaverified] easier, but it is acceptable since we are demonstrating Agent Spec’s functional effectiveness instead of competing on the benchmark leaderboard..
 We also include the performance of a single LLM call as a baseline.
@@
 The first assistant is a ReAct-style agent equipped with a tool to validate that a query can be executed, and a "thinking" tool to give opportunities to the agent to think between tool calls.
-The second assistant is a flow-based assistant composed of a planning step, a step to generate the SQL query and a final post-processing step to self-reflect on the generated query.
+The second assistant is a flow-based assistant  [^5] Details about the flow is provided in Appendix D.2.
+composed of a planning step, a step to generate the SQL query and a final post-processing step to self-reflect on the generated query.
 We also include the performance of a single LLM call as a baseline.
@@
 CrewAI 51.5 / 22.8 $56.1\pm 25.5$
-4 Contributors and Acknowledgements
-Please cite this work as “Open Agent Specification (2025)”.
-Yassine Benajiba, Cesare Bernardis, Vladislav Blinov, Paul Cayet, Hassan Chafi, Abderrahim Fathan, Louis Faucon, Damien Hilloulin, Sungpack Hong, Ingo Kossyk, Rhicheek Patra, Sujith Ravi, Jonas Schweizer, Jyotika Singh, Shailender Singh, Xuelin Situ, Weiyi Sun, Kartik Talamadupula, Jerry Xu, Ying Xu.
-[^2] Author list is sorted alphabetically.
-We are also thankful to our Oracle colleagues who engaged in active discussions during the design of Agent Spec and provided valuable feedback that helped us refine its specifications.
+6 Conclusion & Potential Enhancements
+This paper introduced Agent Spec, a declarative, framework-agnostic specification for defining AI agents and workflows.
+By decoupling design from execution, it enables define-once, run-anywhere portability and interoperability across agentic frameworks.
+We formalized its core components and semantics, and released toolsets: a Python SDK (PyAgentSpec), a reference runtime (WayFlow), and adapters for frameworks (e.g., LangGraph and AutoGen).
+We empirically show that Agent Spec serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across frameworks.
+We outline potential extensions to Agent Spec’s reusability, interoperability, and developer experience in Appendix E.
+Planned enhancements include compiler-style optimizations such as fusing consecutive LLMNodes and using structured generation to reduce redundant calls.
+Looking ahead, we envision Agent Spec as a foundation for standardizing agent design & evaluation.
+Appendix A Simple Agentic Flow in PyAgentSpec
+The following example illustrates the definition of a simple agentic Flow in PyAgentSpec that prompts an LLM with a user-provided input and returns the model’s generated output.
+from pyagentspec.property import Property
+from pyagentspec.flows.flow import Flow
+from pyagentspec.flows.edges import ControlFlowEdge, DataFlowEdge
+from pyagentspec.flows.nodes import LlmNode, StartNode, EndNode
+from pyagentspec.llms import VllmConfig
+llm_config = VllmConfig(
+name="<Your Model Name Here>",
+url="<url.of.your.llm.deployment:port>",
+model_id="<provider/model-identifier>",
+prompt_property = Property(
+json_schema={"title": "prompt", "type": "string"}
+llm_output_property = Property(
+json_schema={"title": "llm_output", "type": "string"}
+start_node = StartNode(name="start", inputs=[prompt_property])
+end_node = EndNode(name="end", outputs=[llm_output_property])
+llm_node = LlmNode(
+name="simple llm node",
+llm_config=llm_config,
+prompt_template="{{prompt}}",
+inputs=[prompt_property],
+outputs=[llm_output_property],
+flow = Flow(
+name="Simple prompting flow",
+start_node=start_node,
+nodes=[start_node, llm_node, end_node],
+control_flow_connections=[
+ControlFlowEdge(name="start_to_llm", from_node=start_node, to_node=llm_node),
+ControlFlowEdge(name="llm_to_end", from_node=llm_node, to_node=end_node),
+data_flow_connections=[
+DataFlowEdge(
+name="prompt_edge",
+source_node=start_node,
+source_output="prompt",
+destination_node=llm_node,
+destination_input="prompt",
+DataFlowEdge(
+name="llm_output_edge",
+source_node=llm_node,
+source_output="llm_output",
+destination_node=end_node,
+destination_input="llm_output"
+Listing 1: Simple Agentic Flow in PyAgentSpec
+To export the Flow that was just created in JSON format, use the following code.
+serializer = AgentSpecSerializer()
+serialized_flow = serializer.to_json(flow)
+Listing 2: Export flow to JSON
+Appendix B Appendix - Implementation Details
+B.1 Agent Spec SDKs
+Agent Spec SDKs are expected to provide two core functionalities:
+  • Construction of Agent Spec component abstractions through the implementation of the corresponding interfaces, in full conformity with the Agent Spec standard.
+  • Import and export of these abstractions to and from their serialized representations in JSON format.
+As part of the Agent Spec initiative, we provide a reference implementation in Python, called PyAgentSpec [pyagentspec].
+PyAgentSpec enables developers to construct Agent Spec-compliant agents directly in Python.
+Users can define their own assistants by instantiating component classes that mirror the interfaces and behaviors prescribed by Agent Spec, and subsequently export these definitions to JSON.
+B.2 Agent Spec SDK Plugins
+A plugin can extend the standard set of Agent Spec components by defining additional component types and specifying the corresponding serialization and deserialization logic required to handle them.
+These plugins enable support for components not yet included in the current version of Agent Spec, such as those related to memory.
+The abstract interface for implementing a serialization plugin is shown below; the deserialization interface follows the same design principles.
+from abc import ABC, abstractmethod
+from typing import Any, Dict, List
+from pyagentspec.component import Component
+from pyagentspec.serialization.serializationcontext import SerializationContext
+class ComponentSerializationPlugin(ABC):
+"""Base class for Component serialization plugins."""
+@property
+@abstractmethod
+def plugin_name(self) -> str:
+"""Return the plugin name."""
+pass
+@property
+@abstractmethod
+def plugin_version(self) -> str:
+"""Return the plugin version."""
+pass
+@abstractmethod
+def supported_component_types(self) -> List[str]:
+"""Indicate what component types the plugin supports."""
+pass
+@abstractmethod
+def serialize(
+self, component: Component, serialization_context: SerializationContext
+) -> Dict[str, Any]:
+"""Serialize a component that the plugin should support."""
+pass
+Listing 3: Interface for serializing an Agent Spec plugin
+B.3 Agent Spec Runtime Adapters
+There are multiple ways to implement runtimes for Agent Spec.
+Ideally, a runtime corresponds to an agentic framework that natively supports Agent Spec, i.e., it can directly read an Agent Spec JSON specification and instantiate an equivalent executable component.
+As a reference example, we provide a runtime implementation called WayFlow [^6] https://github.com/oracle/wayflow, a powerful and intuitive Python library for constructing sophisticated AI-powered assistants.
+WayFlow includes a library of prebuilt plan steps to streamline assistant creation, promotes component reusability, and supports rapid prototyping.
+The framework implements both serialization and deserialization interfaces, enabling the import and export of Agent components that conform to Agent Spec.
+It serves as the reference runtime for Agent Spec, offering native support for all Agent Spec Agents and Flows.
+In cases where an agentic framework does not natively support Agent Spec, a translation layer that transforms the Agent Specifications into a framework-specific representation could be adopted.
+This layer, which we call Agent Spec Runtime Adapter, loads an Agent Specification and converts it into an equivalent representation compatible with the target agentic framework, that can be finally executed.
+We provide runtime adapters [^7] https://github.com/oracle/agent-spec/tree/main/adapters for some of the most popular agentic frameworks, including, but not limited to, LangGraph and Microsoft AutoGen.
+These adapters leverage the Agent Spec Python SDK to load specifications and translate Agent Spec components into their respective framework-specific equivalents.
+Appendix C Appendix - Use Cases & Applications
+A common use case for Agent Spec is the portability of agent-based assistants across different frameworks.
+Consider, for instance, a team that has implemented agent-based workflows using AutoGen.
+As business requirements and process needs evolve (or framework is deprecated [autogenmigration]), the team evaluates alternative frameworks.
+However, due to AutoGen’s specific architecture, migrating these workflows to another agentic framework can be complex and typically entails substantial rewriting overhead.
+In such cases, Agent Spec enables the development of agentic systems, tailored to specific business units, using a declarative representation.
+This representation can then be converted into framework-specific formats or executed flexibly across multiple frameworks via appropriate runtime adapters.
+This approach facilitates direct comparison of task performance and execution costs across frameworks.
+Consequently, employing Agent Spec throughout the agent design and deployment lifecycle significantly reduces complexity, implementation effort, maintenance cost, and potential inconsistencies.
+Another key use case involves the development and deployment of Agents across heterogeneous compute environments, such as cloud platforms and on-premises data centers.
+In cloud-based systems, Agents might run in virtualized or serverless configurations as part of a scalable architecture, whereas in on-premises contexts, they might operate within traditional application servers.
+Agent Spec’s framework and runtime-agnostic design enables development of the Agent in a local development environment, followed by flexible deployment across either of these scenarios.
+An illustrative example of this interoperability is available at https://oracle.github.io/agent-spec/howtoguides/howto_execute_agentspec_across_frameworks.html which demonstrates how an agent defined with Agent Spec can be executed across different frameworks.
+Appendix D Appendix - Agentic Flows
+D.1 RAG Agentic Flow
+The Flow-based Agentic RAG used for SimpleQA is an advanced question-answering system that adapts its reasoning strategy to the complexity of each query.
+The assistant begins with a basic Retrieval-Augmented Generation (RAG) pipeline for straightforward questions and routes to iterative, modular reasoning involving question decomposition, search, result deduplication, and answer synthesis for more complex or open-ended queries.
+Initially, the user’s query is embedded and sent to a web search tool, which retrieves a set of relevant results.
+These results are synthesized into a preliminary answer, which is judged by an LLM to assess if it is likely sufficient to answer the initial user question.
+If the system determines that this initial response adequately answers the user’s question, it is returned immediately, minimizing latency.
+If the response is missing, incomplete or lacks confidence, the system automatically transitions to a more sophisticated, iterative research mode.
+This pipeline begins by constructing a comprehensive context from the user’s original question, any prior sub-questions, retrieved information, and feedback from earlier cycles.
+It then decomposes the complex query into focused sub-questions, which helps target web search and filter out irrelevant details.
+For each sub-question, the system retrieves new results, deduplicates overlapping content, and reassesses previously gathered evidence.
+The assistant then synthesizes a detailed answer and an evaluation component determines whether the answer is now comprehensive; if so, it finalizes and returns the response.
+Otherwise, the refinement loop continues until the assistant achieves a confident response or a maximum number of iterations is reached.
+D.2 NL2SQL Agentic Flow
+This flow incorporates a structured sequence of agentic components to improve reasoning and reliability.
+It begins with an LLM planning node that interprets the user’s query and formulates a high-level strategy.
+The plan is then executed by a ReAct-style agent equipped with external tools for retrieval, computation, or other actions.
+Finally, a post-processing LLM node reviews and refines the agent’s output to verify correctness, check for specified errors, and format the final response before returning it to the user.
+This design was introduced because vanilla ReAct agents based on GPT-4.1 often fail to autonomously generate explicit plans or to perform final verification steps, leading to inconsistent or incomplete results.
+Appendix E Appendix - Potential Enhancements
+E.1 Areas of Potential Extension
+Several extensions could broaden Agent Spec’s applicability to common agentic patterns and enterprise-scale requirements while maintaining its declarative and framework-agnostic design principles.
+Key areas include:
+  • Memory components: Representations for short- and long-term memory, including context persistence and retrieval strategies.
+  • Planning modules: Declarative definitions of planner components for hierarchical or goal-oriented task decomposition.
+  • Datastore interfaces: Native integration with datastores.
+  • Remote agents: Support for defining and communicating with remote agents (e.g., using A2A).
+Each of these enhancements would continue to adhere to the same design principles of clear semantics, explicit I/O, and declarative configuration, while staying framework-agnostic.
+E.2 Runtime Coverage and Conformance
+Expanding runtime coverage remains a core objective.
+We encourage community-led development of runtime adapters for additional agentic frameworks, along with contributions to an extended conformance test suite.
+A more robust test suite will help ensure that identical Agent Spec configurations yield consistent behaviors across diverse runtime environments, thereby improving reproducibility, reliability, and interoperability.
+E.3 Improving User Experience
+While the existing SDKs simplify authoring and validating Agent Specifications, future efforts will focus on improving usability and accessibility, particularly for non-programmatic users.
+Potential directions include:
+  • Visual authoring tools: Drag-and-drop interfaces to construct agents and export Agent Spec configurations to be used in the preferred framework via a runtime adapter.
+  • Advanced diagnostics: Enhanced schema validation, error explanations, and debugging utilities within SDKs.
+These efforts aim to make Agent Spec more accessible for practitioners, educators, and enterprise developers alike.